### v4

# Positional Bias in Long-Document Ranking: Impact, Assessment, and Mitigation

Abstract
We tested over 20 Transformer models for ranking long documents (including recent LongP models trained with FlashAttention and RankGPT models “powered” by OpenAI and Anthropic cloud APIs). We compared them with the simple FirstP baseline, which applied the same model to truncated input (up to 512 tokens). On MS MARCO, TREC DL, and Robust04 no long-document model outperformed FirstP by more than 5% (on average). We hypothesized that this lack of improvement is not due to inherent model limitations, but due to benchmark positional bias (most relevant passages tend to occur early in documents), which is known to exist in MS MARCO. To confirm this, we analyzed positional relevance distributions across four long-document corpora (with six query sets) and observed the same early-position bias. Surprisingly, we also found bias in six BEIR collections, which are typically categorized as short-document datasets. We then introduced a new diagnostic dataset, MS MARCO FarRelevant, where relevant spans were deliberately placed beyond the first 512 tokens. On this dataset, many long-context models—including RankGPT—performed at random-baseline level, suggesting overfitting to positional bias. We also experimented with debiasing training data, but with limited success. Our findings (1) highlight the need for careful benchmark design in evaluating long-context models for document ranking, (2) identify model types that are more robust to positional bias, and (3) motivate further work on approaches to debias training data. We release our code and data to support further research.
(a) Robust04
(b) ClueWeb12 (WebTrack)
(c) TREC DL 2019-2021 (combined)
Figure 1: Positional relevance bias for three long-document collections (best viewed in color). We show a distribution of first relevant passage positions (red bars) vs. relevant document lengths (blue bars). Lengths and offsets are measured in the number of subword tokens (BERT-base tokenizer). See more results (including BEIR) in Figures 7 and 8 in Appendix C.1.
1 Introduction
Various advances in Transformer architectures—including sparse attention Zaheer et al. (2020); Beltagy et al. (2020) and FlashAttention Dao et al. (2022)—have motivated a growing interest in long-document ranking and retrieval. However, despite the ability of these models to process substantially more text, on popular retrieval benchmarks the improvements by these models over simpler truncation-based approaches remain surprisingly modest Dai and Callan (2019); Gao and Callan (2022); Boytsov et al. (2022); Coelho et al. (2024). A widely used truncation-based FirstP baseline Dai and Callan (2019)—where models score only the first 512 tokens of each document—often performs competitively or sometimes even better than long-context counterparts (see, e.g., Table 1).
Despite anecdotal knowledge about the presence of this phenomenon in the MS MARCO document-retrieval collection among TREC Deep Learning track participants and in some early reports Hofstätter et al. (2020b, 2021a); Boytsov et al. (2022) the available evidence has been scattered and incomplete. In particular, despite the track’s five-year history, none of the track’s overview papers have mentioned this issue Craswell et al. (2020, 2021a, 2021b, 2022, 2023).
Moreover, it remains unclear whether these limitations stem from model deficiencies or from characteristics of the benchmarks themselves. In this paper, we initially hypothesized that both factors—model robustness and benchmark design—may be responsible for the limited gains achieved by long-document models over FirstP baselines. However, our findings suggest that benchmark design, particularly positional relevance bias, is the dominant factor.
To verify our research hypotheses, we first conducted a large-scale, systematic study of over 20 Transformer-based ranking models Devlin et al. (2019); Vaswani et al. (2017) for long-document retrieval. This was done using three popular document collections—MS MARCO Documents v1/v2 Craswell et al. (2021a) and Robust04 Clarke et al. (2004)—along with diverse query sets (both large and small), several Transformer backbones, and multiple training seeds. In addition to locally trained models, we also assessed a listwise LLM ranker RankGPT Sun et al. (2023) “powered” by OpenAI OpenAI (2023) and Anthropic Anthropic (2024) cloud APIs.
Despite the increased context capacity of long-document models, we found that none of them consistently outperformed their FirstP baselines by more than 5% on average. An ablation experiment showed that limited gains over FirstP were not attributable to the choice of a Transformer backbone model (see Table 5).
Next, we estimated positional relevance bias across five document collections and more than six query sets. As can be seen from Fig. 1, in the vast majority of cases, the first relevant passage occurred within the initial 512 tokens. In contrast, the distribution of relevant passage positions is more uniform, with a pronounced long tail. This confirms that positional bias is substantial not only in MS MARCO, but also in other TREC collections. Surprisingly, we also found bias in several BEIR collections Thakur et al. (2021), which are typically categorized as short-document datasets (a complete set of plots is provided in Fig. 7 and 8 of Appendix § C.4).
Our initial exploration prompted two broad research questions:
  • RQ1: How robust are long-document models to the positional bias of relevant passages?
[…]
To answer these questions, we constructed a new diagnostic synthetic collection MS MARCO FarRelevant where relevant passages were not present among the first 512 tokens. On this dataset, many long-context models—including RankGPT Sun et al. (2023)—failed to generalize and performed at a random baseline level, suggesting overfitting to positional bias. Poor performance of models on MS MARCO FarRelevant prompted another important question RQ3: Can debiasing of training data mitigate model overfitting to positional bias? We addressed this question by evaluating an existing debiasing approach by Hofstätter et al. (2021a).
Our paper makes the following contributions:
  • We re-examined the issue of positional relevance bias, gathered extensive evidence confirming its presence in both long- and short-document datasets, showed that it negatively affects all models including recent LLM-based rankers, and evaluated the robustness of ranking models against this bias using a new diagnostic dataset MS MARCO FarRelevant.
  • Our work highlights the need for careful benchmark design in evaluating long-context models for document ranking, so that benchmarks do not mask the benefits of long-context models, and identify model types that are more robust to positional bias;
  • We performed an extensive reproduction study of over 20 ranking models using established benchmark collections for long-document retrieval and ranking: MS MARCO Documents v1 and v2 Craswell et al. (2021a) and Robust04 Clarke et al. (2004);
  • We experimented with an existing approach to debiasing training data Hofstätter et al. (2021a) and motivated further research in this area.
Our code and data are available. [^1] https://github.com/searchivarius/long_doc_rank_model_analysis_v2/.
2 Related Work
Neural Ranking models have been a widely studied topic in recent years Guo et al. (2019), though the success of early approaches was debated Lin (2019). This changed with the introduction of BERT, a bidirectional encoder-only Transformer model Devlin et al. (2019), which significantly outperformed previous methods in both NLP Devlin et al. (2019) and information retrieval (IR) tasks Nogueira et al. (2019); Craswell et al. (2021a).
Several Transformer-based models, such as ELECTRA Clark et al. (2020) and DEBERTA He et al. (2021), have improved upon BERT through different training strategies and datasets. However, due to their architectural similarities, we—following Lin et al. (2021)—refer to these collectively as BERT models.
Despite their strong performance, neural models are vulnerable to distribution shifts, often relying on superficial features and exhibiting various biases. They do not consistently outperform BM25 on out-of-domain data Mokrii et al. (2021); Thakur et al. (2021), can be misled by minor text modifications and distractor sentences MacAvaney et al. (2022), or reformulated queries Penha et al. (2022). They also struggle to effectively utilize information located in the middle of long input contexts, in particular for retrieval-augmented generation Liu et al. (2024). Recent work showed that this can be an intrinsic architectural artifact such that the model gives disproportionately high attention to tokens at the beginning or end of the input context, regardless of their actual relevance, which can be partially mitigated by re-calibration Hsieh et al. (2024). However, An et al. (2024) argue that architectural biases can be exacerbated by insufficient explicit supervision during long-context training, which causes models to assume that important information is concentrated near context-window edges.
A related but distinct issue—positional relevance bias—which this study attributes primarily to the characteristics of supervised training data, has been identified in information-retrieval settings, particularly within the MS MARCO document-retrieval collection Hofstätter et al. (2020b); Boytsov et al. (2022); Coelho et al. (2024). Some studies have reported the strong performance of FirstP baselines on long-document retrieval collections, interpreting this as evidence of benchmark-induced positional bias Boytsov et al. (2022); Zhu et al. (2024); Rau et al. (2024). Note, however, that strong performance of FirstP baseline is only indirect evidence, potentially resulting from implementation bugs, suboptimal training methods, or model-inherent biases.
The earliest study on this topic obtained direct evidence of bias only for a small set of TREC DL queries, without studying how bias affects model performance Hofstätter et al. (2020b). Boytsov et al. (2022) tested a variety of long-document ranking models (on MS MARCO and Robust04 datasets) and found them to be only marginally better than the FirstP baseline, which they attributed to positional relevance bias. Coelho et al. (2024) found that two embedding models trained on MS MARCO “dwell” in the beginning and are less effective when relevant information is present elsewhere in a document. These studies, however, (1) had limited evidence that bias was attributable to data, (2) did not evaluate key modern long-context models, and (3) did not assess a debiasing strategy.
Because positional relevance bias can “obscure” benefits of long-document ranking models, Rau et al. (2024) proposed to compare these models with RandP baselines, which score a randomly selected passage. However, we believe this approach is problematic for two reasons. First, since a RandP model often fails to score an entire relevant passage, it artificially underestimates a model’s accuracy, making comparisons with RandP unfair. Second, this approach does not address the core problem of biased benchmarks, which still allow models to exploit the shortcut of focusing only on the document’s initial portion.
Interestingly, Rau et al. (2024) argue that positional bias is strong only in MS MARCO, but not in Robust04. We believe this is because they only compared FirstP and MaxP models in a zero-shot setting (trained on MS MARCO and tested on Robust04), whereas we fine-tuned all models on Robust04 as well. The relatively strong performance of MaxP compared to FirstP in their study may be due to the generally stronger zero-shot performance of MaxP models. We observe similarly small gains over FirstP for both MS MARCO and Robust04, and our passage-relevance analysis confirms that they have comparable levels of positional bias. Thus, despite their widespread use, Robust04 and MS MARCO are not particularly useful for benchmarking of long-document models.
To address the issue with existing benchmarks, Zhu et al. (2024) proposed a LongEmbed benchmark with two synthetic tasks where relevant mini-passages were scattered uniformly across documents with lengths varying from 256 to 32768 tokens. However, as discussed in §C.3, these synthetic sets are quite unnatural and lack diversity. Furthermore, Zhu et al. (2024) provide only small query sets and no in-domain training data, making it difficult to assess the upper performance bound that models can achieve on this dataset. As another important limitation Zhu et al. (2024), only explored training-free extensions of positional encoding and did not investigate methods to debias training data. In contrast, Hofstätter et al. (2021a) proposed to debias training data using a simple yet effective approach. However, they did not evaluate it on challenging long-document datasets.
In this paper, we aim to provide stronger evidence through collecting more comprehensive statistics of relevant passages in documents, experimenting with a diverse set of models, and applying the diagnostic synthetic set, which also demonstrates models’ abilities to adapt to different positional biases.
Due to the quadratic complexity of the Transformer’s attention mechanism Vaswani et al. (2017); Bahdanau et al. (2015), early Transformer models restricted input length to a maximum of 512 (subword) tokens. Until around 2022, two main strategies were used to process long documents: (1) localizing attention and (2) splitting documents into smaller, independently processed chunks. Attention-localization methods apply a limited-span (sliding window) attention and selective global attention. Given the vast number of such approaches (see Tay et al. 2020), evaluating all of them is impractical. Therefore, we focus on two popular models: Longformer Beltagy et al. (2020) and BigBird Zaheer et al. (2020). More recently, it has also become feasible to train long-context models with an IO-efficient FlashAttention algorithm without sparsifying attention Dao et al. (2022).
In summary, methods to tackle longer documents are divided into LongP methods—where longer document are “natively” supported and SplitP methods—where a longer document cannot be processed as a whole and needs to be processed in chunks. The results of each chunk are aggregated together using various aggregation techniques, including computation of a maximum or a weighted-sum prediction score Yilmaz et al. (2019); Dai and Callan (2019); MacAvaney et al. (2019). This includes MaxP Dai and Callan (2019), AvgP, SumP MacAvaney et al. (2019), as well as PARADE Avg and PARADE Max models MacAvaney et al. (2019). MaxP is an important baseline that computes relevance scores for each chunk independently and takes their maximum.
Some SplitP approaches aggregate using simple neural networks. This includes all CEDR MacAvaney et al. (2019) models, the Neural Model 1 Boytsov and Kolter (2021), and the PARADE Attention model Li et al. (2024). In contrast, PARADE Transformer Li et al. (2024) models’ aggregator network is an additional Transformer model. Due to space constraints, a detailed description of document-splitting (SplitP) approaches is provided in the Appendix § D.
The recent success of decoder-only models—commonly known as LLMs—has led to a new generation of cross-encoding LongP models that natively support longer contexts. First, a pre-trained cross-encoding decoder-only model can be directly fine-tuned on a ranking or embedding task Ma et al. (2023). Second, the RankGPT approach Sun et al. (2023) formulates document ranking as a generation task: The model is prompted with a list of documents and an instruction to generate their ranking—an ability made possible through instruction-tuning and/or alignment Wei et al. (2022); Ouyang et al. (2022). When the combined length of concatenated documents exceeds the input context size, RankGPT employs an overlapping sliding window strategy, followed by aggregation of the results.
Retriever / Ranker MS MARCO TREC DL Robust04 Avg. gain
dev (2019-2021) title description over FirstP
MRR NDCG@10 NDCG@20
BM25 0.274 0.545 0.428 0.402
retriever (if different from BM25) 0.312 0.629
FirstP (BERT) 0.394 0.632 0.475 0.527
FirstP (Longformer) 0.404 0.643 0.483 0.540
FirstP (ELECTRA) 0.417 0.662 0.492 0.552
FirstP (DEBERTA) 0.415 0.672 0.534 0.596
FirstP (Big-Bird) 0.408 0.656 0.507 0.560
FirstP (JINA) 0.422 0.654 0.488 0.532
FirstP (MOSAIC) 0.423 0.643 0.453 0.538
FirstP (TinyLLAMA) 0.395 0.615 0.431 0.473
FirstP (E5-4K) zero-shot 0.380 0.641 0.438 0.429
FirstP RankGPT (GPT-4o-mini) 0.708 0.562
AvgP 0.389 $(-1.3\%)$ 0.642 $(+1.5\%)$ 0.478 $(+0.5\%)$ 0.531 $(+0.9\%)$ $0.4\%$
MaxP 0.392 $(-0.4\%)$ 0.644a $(+1.9\%)$ 0.488a $(+2.6\%)$ 0.544a $(+3.3\%)$ $1.9\%$
MaxP (ELECTRA) 0.414 $(-0.6\%)$ 0.659 $(-0.5\%)$ 0.502 $(+2.0\%)$ 0.563 $(+2.1\%)$ $0.8\%$
MaxP (DEBERTA) 0.402a $(-3.2\%)$ 0.671 $(-0.1\%)$ 0.535 $(+0.2\%)$ 0.609a $(+2.2\%)$ -$0.2\%$
SumP 0.390 $(-1.0\%)$ 0.639 $(+1.0\%)$ 0.486 $(+2.2\%)$ 0.538 $(+2.1\%)$ $1.1\%$
CEDR-DRMM 0.385a $(-2.3\%)$ 0.629 $(-0.5\%)$ 0.466 $(-2.0\%)$ 0.533 $(+1.3\%)$ -$0.9\%$
CEDR-KNRM 0.379a $(-3.8\%)$ 0.630 $(-0.3\%)$ 0.483 $(+1.7\%)$ 0.535 $(+1.7\%)$ -$0.2\%$
CEDR-PACRR 0.395 $(+0.3\%)$ 0.643a $(+1.6\%)$ 0.496a $(+4.3\%)$ 0.549a $(+4.2\%)$ $2.6\%$
Neural Model1 0.398 $(+0.9\%)$ 0.650a $(+2.8\%)$ 0.484 $(+1.8\%)$ 0.537 $(+1.9\%)$ $1.8\%$
PARADE Attn 0.416a $(+5.5\%)$ 0.652a $(+3.1\%)$ 0.503a $(+5.7\%)$ 0.556a $(+5.6\%)$ +5.0%
PARADE Attn (ELECTRA) 0.431a $(+3.3\%)$ 0.680a $(+2.7\%)$ 0.523a $(+6.4\%)$ 0.581a $(+5.3\%)$ $4.4\%$
PARADE Attn (DEBERTA) 0.422a $(+1.6\%)$ 0.688a $(+2.4\%)$ 0.549a $(+2.9\%)$ 0.615a $(+3.2\%)$ $2.5\%$
PARADE Avg 0.392 $(-0.6\%)$ 0.646a $(+2.1\%)$ 0.483 $(+1.5\%)$ 0.534 $(+1.5\%)$ $1.1\%$
PARADE Max 0.405a $(+2.7\%)$ 0.655a $(+3.5\%)$ 0.489a $(+2.8\%)$ 0.548a $(+4.0\%)$ $3.3\%$
PARADE Transf-RAND-L2 0.419a $(+6.3\%)$ 0.655a $(+3.6\%)$ 0.488a $(+2.8\%)$ 0.548a $(+4.1\%)$ $4.2\%$
PARADE Transf-RAND-L2 (ELECTRA) 0.433a $(+3.9\%)$ 0.670 $(+1.2\%)$ 0.523a $(+6.3\%)$ 0.574a $(+3.9\%)$ $3.8\%$
PARADE Transf-PRETR-L6 0.402a $(+1.9\%)$ 0.643 $(+1.6\%)$ 0.494a $(+4.0\%)$ 0.554a $(+5.1\%)$ $3.2\%$
LongP (Longformer) 0.412a $(+1.9\%)$ 0.668a $(+3.9\%)$ 0.500a $(+3.6\%)$ 0.568a $(+5.1\%)$ $3.6\%$
LongP (Big-Bird) 0.397a $(-2.9\%)$ 0.651 $(-0.7\%)$ 0.452a $(-10.9\%)$ 0.477a $(-14.9\%)$ -$7.3\%$
LongP (JINA) 0.416a $(-1.5\%)$ 0.665a $(+1.7\%)$ 0.503a $(+2.9\%)$ 0.558a $(+4.9\%)$ $2.0\%$
LongP (MOSAIC) 0.421 $(-0.4\%)$ 0.664a $(+3.3\%)$ 0.456 $(+0.6\%)$ 0.570a $(+6.0\%)$ $2.4\%$
LongP (TinyLLAMA) 0.402a $(+1.7\%)$ 0.608 $(-1.1\%)$ 0.452a $(+4.8\%)$ 0.505a $(+6.7\%)$ $3.0\%$
LongP (E5-4K) zero-shot 0.353a $(-7.1\%)$ 0.649 $(+1.3\%)$ 0.439 $(+0.1\%)$ 0.434 $(+1.1\%)$ -$1.1\%$
LongP RankGPT (GPT-4o-mini) 0.706 $(-0.3\%)$ 0.562 $(+0.0\%)$ -$0.1\%$
  • In each column we show a relative gain with respect model’s respective FirstP baseline: The last column shows the average relative gain over FirstP baselines. Best numbers are in bold: Results are averaged over three seeds. Unless specified explicitly, the backbone is BERT-base. Statistical significant differences with respect to this baseline are denoted using the superscript a. $p$-value threshold is 0.01 for an MS MARCO development collection and 0.05 otherwise.
Table 1: Comparison between long-document models and respective FirstP (truncation) baselines. Results for MS MARCO, TREC DL, and Robust04.
3 Experiments
3.1 Data
Our primary datasets, used for both training and evaluation, consist of several realistic collections (along with their respective query sets) and synthetic data. All datasets are in English. Document and query statistics are provided in Appendix § C.1; see Tables 7 and 8.
The realistic datasets include three long-document collections: MS MARCO Documents v1 and v2 Bajaj et al. (2016); Craswell et al. (2020, 2021b), Robust04 Voorhees (2004), and several short-document collections: MS MARCO Passages (v1) Bajaj et al. (2016); Craswell et al. (2020, 2021b), and seven BEIR datasets (see § C.1). Several BEIR datasets and the MS MARCO Documents collections include document titles, which we prepend to the main document text. Following Hofstätter et al. (2021a), we created a debiased version of MS MARCO by randomly splitting documents at word boundaries and concatenating reordered segments (see Algorithm 1 in § A.1).
Our synthetic data consists of two subsets from LongEmbed Zhu et al. (2024) and our newly created MS MARCO FarRelevant collection. All these can be considered variants of the needle-in-the-haystack test, where an informational “nugget” is randomly embedded within unrelated text Saad-Falcon et al. (2024); Zhu et al. (2024); Liu et al. (2024). We use two LongEmbed subsets: Needle and Passkey. Each subset has 800 question-document pairs with document lengths varying from approximately 256 to 32768 tokens.
MS MARCO FarRelevant was created by randomly mixing relevant and non-relevant passages from the MS MARCO Passage collection Craswell et al. (2020) in such a way that (1) each document contains exactly one relevant passage, (2) this passage does not start before token 512, and (3) less than 1% of documents have more than 1500 tokens (see an algorithm in Appendix § C.2). It has about 0.5 million documents with an average length of 1.1K tokens. Due to the MS MARCO datasets’ non-commercial license, MS MARCO FarRelevant has the same licensing restriction. In Appendix § C.3, we present dataset examples and argue that—while all these collections share the limitation of not resembling natural documents—MS MARCO FarRelevant offers greater diversity and serves as a more suitable benchmark for evaluating text retrieval systems.
Robust04 is another relatively small dataset containing 0.5 million documents, comprising a mix of news articles and government records, some of which are quite lengthy. However, it includes only a limited number of queries (250), making it a challenging benchmark for training models in low-data scenarios. Each query has a title and a description. The title expresses a concise information need, while the description provides a more detailed request, often written in proper English prose. We use Robust04 in a cross-validation setting with folds created by Huston and Croft (2014) and provided via IR-datasets MacAvaney et al. (2021). [^2] In this setting we do not train Robust04 models from scratch, but instead fine-tune models trained on the MS MARCO Documents collection.
MS MARCO v1 was created from the MS MARCO reading comprehension dataset Bajaj et al. (2016) and consists of two related collections: MS MARCO Passages and MS MARCO Documents. MS MARCO v1 comes with large query sets, which is particularly useful for training and testing models in the big-data regime. These query sets include question-like queries selected from the Bing search engine logs Craswell et al. (2021b). Note that queries are not necessarily proper English questions, e.g., “lyme disease symptoms mood”, but they are answerable by a short passage retrieved from a set of about 3.6M Web documents Bajaj et al. (2016). MS MARCO v1 test sets were created in two stages. First, relevance judgments were created for the passage variant of the dataset. Then, document-level relevance labels were created by transferring passage-level relevance to original documents from which passages were extracted.
The MS MARCO v2 collection was created for the TREC 2021 Deep Learning (DL) track Craswell et al. (2021b). It is an expanded version of MS MARCO v1 and incorporates a subset of sparse relevance judgments from MS MARCO v1. In the training set, newly added documents lack both positive and negative judgments, introducing a bias where many relevant documents are mistakenly considered non-relevant. As a result, we do not train on v2 data and only use it for testing.
Relevance labels in the training and development sets are “sparse”: There is about one positive example per query without explicit negatives. In addition to sparse relevance judgments—separated into training and development subsets—there is a small number of (about 150) test queries from the TREC Deep Learning (DL) track.
3.2 Setup
We focus on cross-encoding rankers, which process queries concatenated with documents Nogueira and Cho (2019). This includes various SplitP and LongP models discussed in § 2 and in the Appendix § D. As a reference point, we also tested a bi-encoder embedding E5-4K model, which has strong performance on the LongEmbed benchmark with context sizes under 4K tokens Zhu et al. (2024). E5-4K was tested only in zero-shot mode (without fine-tuning)—as a ranking model. Except for LongEmbed subsets Needle and Passkey, a ranker is applied to a top-$k$ set produced by a retriever. Each LongEmbed subset contains only 800 documents, and we re-rank them all without using a first-stage retriever.
Nearly all rankers are based on BERT models (bidirectional encoder-only Transformer) with 100-200M parameters (see Table 11). Additionally, we evaluated two types of LLM rankers: (1) a fine-tuned TinyLLAMA model, which delivers strong performance relative to its compact size Zhang et al. (2024) and (2) generative black-box LLMs. For (2), we used OpenAI’s GPT-4o-mini OpenAI (2023) and Anthropic’s Claude Haiku-3 Anthropic (2024), both of which support at least a 128K-token input context.
Unless explicitly specified, the backbone Transformer model for SplitP methods is BERT-base Devlin et al. (2019). Although using other backbones such as ELECTRA Clark et al. (2020) and DEBERTA He et al. (2021) can improve overall accuracy, we observe larger gains compared to the FirstP baseline when we use BERT-base (see § B.1 in the Appendix).
Except for E5 Zhu et al. (2024) and RankGPT Sun et al. (2023), which were evaluated only in zero-shot mode, we trained each model using three seeds. Due to the high evaluation cost (more than \$1 per 1000 query-document pairs) [^3] https://openai.com/api/pricing/, we also did not test RankGPT on some query sets, in particular, we excluded the large MS MARCO development set.
We measure effectiveness of the models using the mean reciprocal rank (MRR), the normalized discounted cumulative gain at rank $k$ (NDCG@k) Järvelin and Kekäläinen (2002), precision at rank $k$ (P@k), and the mean average precision (MAP). To assess statistical significance, we averaged per-query metric values across three seeds.
Figure 2: Zero-shot vs. fine-tuned performance on MS MARCO FarRelevant. This figure shows results for a representative set of models.
3.3 Results
Realistic Data.
Our main experimental results for MS MARCO, TREC DL 2019-2021, and Robust04 are presented in Table 1. There we show only a single aggregate number for all TREC DL query sets, which is obtained by combining all the queries and respective relevance judgments (i.e., we post an overall average rather than an average over the mean values for 2019, 2020, and 2021). More detailed results, including Anthropic-based RankGPT, are available in Appendix B, specifically in Tables 9 and 10. Efficiency evaluation is presented in § B.2.2 (see Fig. 3). In § B.1 we also show that we can match or outperform key prior results, which—we believe—boosts the trustworthiness of our experiments.
We abbreviate names of some PARADE models: Note that PARADE Attn denotes a PARADE Attention model. The PARADE Transf or P. Transf prefix denotes PARADE Transformer models where an aggregator Transformer can be either trained from scratch (Transf-RAND-L2) or initialized with a pretrained model (Transf-PRETR-L6). L2 and L6 denote the number of aggregating layers (two and six, respectively). [^4] Note, however, that Transf-PRETR-L2 has only four attention heads.
From Table 1 and Fig. 3 we learn that the maximum average gain relative to respective FirstP baselines is only 5% (when measured using MRR or NDCG@k). Gains are much smaller for a number of models, which sometimes even underperform their FirstP baselines on one or more datasets. In particular, this is true for RankGPT Sun et al. (2023), CEDR-DRMM, CEDR-KNRM MacAvaney et al. (2019), JINA Günther et al. (2023), and MOSAIC Portes et al. (2023).
Ranker MS MARCO TREC DL FarRelevant LongEmbed
dev (2019-2021) zero-shot transf. fine-tuned Needle Passkey
MRR NDCG@10 MRR MRR MRR MRR
BM25 0.274 0.545 0.207 0.207 0.305 0.339
Original MS MARCO training set
FirstP (ELECTRA) 0.417 0.662 0.019 0.089 0.205 0.235
MaxP (ELECTRA) 0.414 0.659 0.328 0.349 0.331 0.338
PARADE Attn (ELECTRA) 0.431 0.680 0.338 0.354 0.270 0.334
PARADE Transf-RAND-L2 (ELECTRA) 0.433 0.670 0.229 0.432 0.321 0.333
CEDR-KNRM 0.379 0.630 0.055 0.382 0.129 0.166
Debiased MS MARCO Hofstätter et al. (2021a)
MaxP (ELECTRA) 0.377a $(-9.1\%)$ 0.665 $(+0.8\%)$ 0.321 $(-2.1\%)$ 0.349 0.316 $(-4.6\%)$ 0.325 $(-3.9\%)$
PARADE Attn (ELECTRA) 0.390a $(-9.4\%)$ 0.653a $(-3.9\%)$ 0.326 $(-3.6\%)$ 0.354 0.251 $(-7.1\%)$ 0.330 $(-1.0\%)$
PARADE Transf-RAND-L2 (ELECTRA) 0.410a $(-5.4\%)$ 0.677 $(+1.0\%)$ 0.328a $(+43.5\%)$ 0.432 0.259a $(-19.4\%)$ 0.331 $(-0.7\%)$
CEDR-KNRM 0.269a $(-29.0\%)$ 0.503a $(-20.2\%)$ 0.202a $(+268.8\%)$ 0.382 0.121 $(-5.8\%)$ 0.181 $(+8.6\%)$
  • Except FirstP we train each model using the original and the debiased MS MARCO. For each model trained on the debiased dataset, we compute a gain (or loss) compared to the same model trained on the original training set. Statistical significance of the respective difference is denoted using the superscript superscript a ($p$-value threshold is 0.05 for TREC DL and 0.01 for other collections). Best numbers are in bold: Results are averaged over three seeds.
Table 2: Debiasing effectiveness. Performance of (selected) rankers trained on original and debiased MS MARCO.
We can also see that the LongP variant of the Longformer model appears to have a relatively strong performance, but so does the FirstP version of Longformer. Thus, we think that a good performance of Longformer on MS MARCO and Robust04 collections can be largely explained by better pretraining compared to the original BERT-base model rather than by its ability to process long contexts. Moreover, FirstP (ELECTRA) and FirstP (DEBERTA) are even more accurate than FirstP (Longformer) and perform comparably well to (or better than) some chunk-and-aggregate document models that use BERT-base as the backbone model. This is a fair comparison aiming to demonstrate that—on a typical test collection—the benefits of long-context models are so small that comparable benefits can be obtained by finding or training a more effective FirstP model. FirstP models are more efficient during inference and can be pretrained on a larger number of tokens for the same cost, so they may perform better (see Fig. 3 in § B.2.2).
Synthetic Data.
Based on our analysis of positions of first relevant passages, we hypothesized that limited benefits of long-context models are not due to inability to process long contexts, but rather due to a positional bias of relevant passages, which tend to be among the first 512 document tokens (see Figure 1 and Figure 7 in Appendix C.4). To support this hypothesis, we carried out two sets of experiments using MS MARCO FarRelevant, where a relevant passage did not start until token 512. We carried out both a zero-shot experiment (evaluation of the model trained on MS MARCO) and a fine-tuning experiment using 50K in-domain queries (from the MS MARCO FarRelevant).
Results for key models are shown in Fig. 2 and more detailed results can be found in Table 4 of the Appendix B. The FirstP models performed roughly at the random-baseline level in both zero-shot and fine-tuning modes (RQ1). Because of this, our main baselines here are Longformer and MaxP models. For models with ELECTRA and DEBERTA backbones, we compare with MaxP (ELECTRA) and MaxP (DEBERTA), respectively. Otherwise, the baseline is MaxP (BERT).
Surprisingly, E5-4K performance is also at a random-baseline level despite its competitive performance on LongEmbed benchmark Zhu et al. (2024), MS MARCO, and Robust04 (see Table 1). Both GPT-4o-mini and Claude Haiku-3 RankGPT perform at the random-baseline level as well! As a sanity check, and to verify if more accurate and expensive LLMs could do better, we assessed performance of GPT-4o for a sample of 100 queries. The respective RankGPT Sun et al. (2023) ranker was still not better than a random baseline (RQ1).
Simple aggregation models including MaxP and PARADE Attention had good zero-shot accuracy, but benefited little from fine-tuning on MS MARCO FarRelevant (RQ1).
In contrast, other long-document models had poor zero-shot performance (sometimes at a random baseline level), but outstripped respective MaxP baselines by as much as 9.2%-27.7% after fine-tuning, see Table 4 and Fig. 2 (RQ1 and RQ2).
With the exception of RankGPT Sun et al. (2023) on TREC DL 2019-2021, PARADE Transformer models were more effective than other models on standard collections, but their advantage was small (a few %). In contrast, on MS MARCO FarRelevant, PARADE Transformer (ELECTRA) outperformed the next competitor Longformer by 8% and PARADE Max (ELECTRA)—an early chunk-and-aggregate approach—by as much as 23.8%, thus showcasing substantial advances in long-document ranking modeling and answering RQ2.
Bias Mitigation.
To address RQ3, we trained four representative models on the debiased MS MARCO and tested them on MS MARCO FarRelevant as well as Needle/Passkey subsets of LongEmbed Zhu et al. (2024).555(see Appendix § A.1.3 for model selection rationale) We also tested four models fine-tuned on MS MARCO FarRelevant on TREC DL 2019-2021 query sets. Due to the substantial NDCG@10 drop (0.1–0.15) observed for PARADE Transformer and CEDR-KNRM, we concluded that fine-tuning on purely synthetic data is not viable and did not pursue it further.
According to Table 2, debiasing improved the performance of CEDR-KNRM and PARADE Transformer on MS MARCO FarRelevant. Yet, it mostly caused performance degradation on the original MS MARCO dataset and on LongEmbed subsets. It did not benefit the MaxP and PARADE Attention models, which were the most robust to positional bias. We believe this result—along with the strong performance of all models after fine-tuning on MS MARCO FarRelevant—supports the conclusion that the models are not inherently biased toward the start of the document (or at least this is not a key factor). Instead, the primary cause appears to be positional bias in the training data.
We further tested debiased models on short-document collections: the MS MARCO Passage collection and seven BEIR collections Thakur et al. (2021). According to Table 3 in Appendix A, for three out of four models, debiasing has either a positive or small negative effect (at most 1% degradation). In particular, on the BEIR NQ subset, PARADE Transformer and CEDR-KNRM improved by 2.5% and 11%, respectively.
Despite the positive average improvement, we also observe substantial degradation for several datasets. For example, the effectiveness of CEDR-KNRM decreases by 12.4% and 4.6% on SciFact and SciDocs, respectively. One possible reason is that short-document collections can have positional relevance bias as well (see Figure 8 in Appendix C.1). Thus, training on the original (biased) MS MARCO collection teaches the model to rely on a helpful shortcut—to focus more on the beginning of a document—whereas training on debiased data removes this beneficial behavior.
In summary, debiasing results are promising, but they also suggest that mitigating positional bias remains a challenging problem (RQ3).
Key Findings.
We observe the following:
  • We confirmed the presence of positional relevance bias in both long- and short-document datasets.
  • Not only did positional relevance bias diminish the benefits of processing longer document contexts, but it also led to model overfitting to the bias and performing poorly in a zero-shot setting when the distribution of relevant passages changes substantially.
  • We found that debiasing training data had mixed success: Although it improved the effectiveness of some models on some long-document ranking tasks without substantial degradation on short-document collections, it degraded performance in several cases—especially for MaxP and PARADE Attention models—which were already intrinsically more robust to positional relevance bias—possibly because many short-document collections also have positional relevance bias.
  • It is also worth highlighting the consistently strong performance of PARADE models on both standard long-document collections and MS MARCO FarRelevant. The best PARADE models substantially outperformed the best LongP models in both zero-shot and fine-tuning settings, although the specific models leading in each setting may differ.
4 Conclusion
In this work, we revisited the problem of positional relevance bias in long-document retrieval and presented extensive evidence of its widespread presence across existing benchmarks, including several BEIR datasets Thakur et al. (2021), which contain relatively short documents. Using both real and synthetic datasets—including our new diagnostic dataset MS MARCO FarRelevant—we evaluated the effectiveness of over 20 ranking models as well as their robustness to positional relevance bias.
Our findings highlight the importance of a benchmark design that does not obscure the benefits of long-context modeling. We identified model families (i.e., PARADE Attention and MaxP) that are more robust to positional bias, and confirmed the strong performance of PARADE models Li et al. (2024), which remain competitive even against recent long-context architectures.
Finally, our debiasing experiments yielded limited and inconsistent gains, motivating further research into more effective mitigation strategies, including combining debiasing with training on well-designed synthetic data.
5 Limitations
Our paper has several limitations related primarily to the choice of datasets, models, and strength of evidence for positional relevance bias.
First of all, our evaluation uses only cross-encoding ranking models. With the exception of E5-4K model, which is used in the zero-shot ranking mode, we do not train or evaluate bi-encoding models (typically used to create query and document embeddings for the first-stage retrieval). We nonetheless believe that—given a large number of proposals for long-document ranking—a reproduction and evaluation of cross-encoding long-document rankers is a sufficiently important topic that alone warrants a publication.
Moreover, we also use cross-encoding rankers as a tool to detect and expose positional relevance bias. In this regard, cross-encoders are easier to train using standard (rather than high-memory) GPUs with a mini-batch size of one and gradient accumulation. They also typically require only one epoch to converge (only a few models need two or three epochs). In contrast, bi-encoders are trained using large batches with in-batch negatives for multiple epochs (e.g., Karpukhin et al. (2020) report using at least 40 epochs).
Second, the bulk of our ranking experiments uses only two English document collections: MS MARCO Documents v1 and v2 Craswell et al. (2021b) and Robust04 Clarke et al. (2004). However, we have to restrict the choice of datasets to make multi-seed evaluations of 20+ models feasible. Thus, to corroborate the widespread presence of positional bias we used two additional popular long-document collections: Gov2 Allan et al. (2008) and ClueWeb12 Collins-Thompson et al. (2013a) as well as seven BEIR datasets Thakur et al. (2021). For the study on the robustness of models to positional bias and its mitigation, we used two additional synthetic collections: MS MARCO FarRelevant and two subsets from LongEmbed Zhu et al. (2024) together with eight short-document collections (see Table 3).
One could argue that the limited improvements over FirstP baselines result from the models’ inability to handle long contexts. To address this concern, we trained and evaluated a diverse set of cross-encoding ranking models, including both split-and-aggregate models and models explicitly designed for long input sequences. Additionally, we assessed cloud-based RankGPT rankers, which have shown strong performance in recent research Sun et al. (2023).
However, we can still test only a limited number of models. In this regard, one might always argue that there are untested architectures that would outperform FirstP baselines by a much larger margin. To demonstrate that selected models can, in principle, benefit from long contexts and decisively outperform simple baselines such as FirstP and even MaxP models, we trained and/or evaluated them on a synthetic collection MS MARCO FarRelevant, which can be seen as a challenging version of a needle-in-the-haystack test.
Admittedly, this is still a limiting experiment, because synthetic collections—with documents composed of randomly selected passages—do not represent complete and coherent documents. In § C.3 we discuss this limitation in detail; nevertheless, we argue that MS MARCO FarRelevant is a more suitable synthetic benchmark for evaluating text retrieval systems compared to LongEmbed subsets Needle and Passkey Zhu et al. (2024).
In summary, we provided three types of evidence of positional bias in relevant passages: strong performance of FirstP models on standard collections, direct estimation of the distribution of relevant passages using substring matching and LLM relevance judges Upadhyay et al. (2024), as well as experimentation with the synthetic collection MS MARCO FarRelevant where the distribution of relevant passages was not skewed towards the beginning of a document. Each experiment provided imperfect or limited evidence on its own, but together they strongly support the existence of positional relevance bias.
While our analysis confirms a strong early-position relevance bias across multiple retrieval benchmarks, we acknowledge that this pattern may not generalize to all domains. For example, prior work has shown that in scientific abstracts, both the first and the last sentences tend to be crucial Ruch et al. (2006). Investigating positional relevance patterns in such domains is an important direction for future work.
In our experiments with Robust04 and MS MARCO, we truncated documents to a maximum of 1431 BERT tokens. However, this constraint did not hinder our ability to address key research questions. As detailed in Appendix § B.2.1, using larger inputs led to only marginal improvements.
Appendix A Experimental Addendum: Training/Evaluation Setup, Ablations, and Detailed Results
A.1 Detailed Training and Evaluation Setup
A.1.1 General Setup
Except for LongEmbed Zhu et al. (2024), we applied a ranker to the output of the first-stage retrieval model, also known as a candidate generator. However, each LongEmbed subset contains only 800 documents, and we re-rank them all without using a first-stage retriever.
Depending on the experiment and the dataset, we used different candidate generators. For MS MARCO v1 (training and development sets) and Robust04, we used a BM25 ranker Robertson (2004). Note that for MS MARCO v1, the ranker was applied to documents expanded using the doc2query approach Nogueira and Lin (2019).
For MS MARCO v2, we used a hybrid retriever where candidate records were first produced using a k-NN search and subsequently re-ranked using a linear combination of BM25 scores and the cosine similarity between query and document embeddings. We also used k-NN search with neural embeddings for TREC DL queries. Embeddings were generated using ANCE Xiong et al. (2021).
Retriever / Ranker zero-shot fine-tuned
transferred
Random shuffling of top-100 0.052 0.052
Retriever (BM25) $^{\phantom{b}} 0.207{}^{\phantom{b}}$
FirstP (BERT) 0.016b 0.090b
FirstP (Longformer) 0.017b 0.091b
FirstP (ELECTRA) 0.019b 0.089b
FirstP (Big-Bird) 0.021b 0.089b
FirstP (JINA) 0.018b 0.088b
FirstP (MOSAIC) 0.018b 0.089b
FirstP (TinyLLAMA) 0.020b 0.079b
FirstP (E5-4K) 0.015ab
AvgP 0.154ab $(-48.1\%)$ 0.365ab $(+11.4\%)$
MaxP 0.297b 0.328b
MaxP (ELECTRA) 0.328b 0.349b
MaxP (DEBERTA) 0.298b 0.332b
SumP 0.211ab $(-28.8\%)$ 0.327b $(-0.4\%)$
CEDR-DRMM 0.157ab $(-47.3\%)$ 0.372ab $(+13.3\%)$
[…]
PARADE Attn 0.300b $(+1.0\%)$ 0.337b $(+2.8\%)$
PARADE Attn (ELECTRA) 0.338b $(+3.3\%)$ 0.354b $(+1.6\%)$
PARADE Attn (DEBERTA) 0.307b $(+3.2\%)$ 0.343b $(+3.4\%)$
PARADE Avg 0.274ab $(-7.6\%)$ 0.322b $(-1.7\%)$
PARADE Max 0.291b $(-2.1\%)$ 0.330b $(+0.6\%)$
[…]
LongP (JINA) 0.069ab $(-76.9\%)$ 0.372ab $(+13.4\%)$
[…]
LongP (E5-4K) 0.057ab $(-80.7\%)$ N/A (zero-shot only)
LongP RankGPT (GPT-4o-mini) 0.043b N/A (zero-shot only)
LongP RankGPT (Claude-3-haiku) 0.051b N/A (zero-shot only)
  • In each column we show a relative gain over models’ respective MaxP baseline. For LongP models, the gain is over MaxP (BERT).
    Statistically significant differences from a respective MaxP baseline are denoted with the superscript a.
    Statistical significant differences with respect to Longformer are denoted with the superscript b ($p$-value <0.01<0.01).
Table 4: Comparison between long-document models and respective FirstP (truncation) baselines. Results on MS MARCO FarRelevant.
Depending on the collection we computed a subset of the following metrics: the mean reciprocal rank (MRR), the normalized discounted cumulative gain at rank $k$ (NDCG@k) Järvelin and Kekäläinen (2002), the mean average precision (MAP), and precision at rank (P@k), $k\in\{10,20\}$. Due to space constraints, we included results with MAP and P@K only in the Appendix (see § B). Note that for test sets with sparse labels (MS MARCO development set and MS MARCO FarRelevant) we computed only MRR.
All experiments were carried out using the FlexNeuART framework Boytsov and Nyberg (2020), which employed Lucene and NMSLIB Boytsov and Naidan (2013) to provide retrieval capabilities. Deep learning support was provided via PyTorch Paszke et al. (2019) and HuggingFace Transformers library Wolf et al. (2019). The instructions to reproduce our key results are publicly available. [^6] https://github.com/searchivarius/long_doc_rank_model_analysis_v2/
A.1.2 Model Training
A ranker was trained to distinguish between positive examples (known relevant documents) and hard negative examples (documents not marked as relevant) sampled from the set of top-$k$ candidates returned by the candidate generator. Based on preliminary experiments, we chose $k=100$ for MS MARCO and MS MARCO FarRelevant. For Robust04 we used $k=1000$.
Each model was trained using three seeds. To compute statistical significance, we averaged query-specific metric values over these seeds. All models except MOSAIC were trained using half-precision. MOSAIC models were trained using full-precision. MOSAIC training was unstable (even though we used the full precision) and often resulted in close-to-zero performance. For this reason we continued training with more seeds until we obtained three models with reasonable performance. This seed selection strategy could potentially have biased (up) MOSAIC results.
Model MS MARCO TREC DL Robust04
dev 2019 2020 2021 title description
MRR NDCG@10 NDCG@20
BM25 0.274 0.548 0.538 0.549 0.428 0.402
Prior work (FirstP, MaxP), Zhang et al. Zhang et al. (2021)
FirstP (BERT) 0.449 0.510
MaxP (BERT) 0.477 $(+6.2\%)$ 0.530 $(+3.9\%)$
MaxP (ELECTRA) 0.523 0.574
Prior work (PARADE) Li et al. Li et al. (2024)
PARADE Attn (ELECTRA) 0.527 0.587
PARADE Max (ELECTRA) 0.679 0.613 0.544 0.602
PARADE Transf-RAND (ELECTRA) 0.650 0.601 0.566 0.613
Our results
FirstP (BERT) 0.394 0.631 0.598 0.660 0.475 0.527
MaxP (BERT) 0.392 $(-0.4\%)$ 0.648 $(+2.6\%)$ 0.615 $(+2.8\%)$ 0.665 $(+0.8\%)$ 0.488a $(+2.6\%)$ 0.544a $(+3.3\%)$
PARADE Attn 0.416a $(+5.5\%)$ 0.647 $(+2.5\%)$ 0.626a $(+4.6\%)$ 0.677 $(+2.5\%)$ 0.503a $(+5.7\%)$ 0.556a $(+5.6\%)$
FirstP (ELECTRA) 0.417 0.652 0.642 0.686 0.492 0.552
MaxP (ELECTRA) 0.414 $(-0.6\%)$ 0.659 $(+1.0\%)$ 0.630 $(-1.9\%)$ 0.683 $(-0.5\%)$ 0.502 $(+2.0\%)$ 0.563 $(+2.1\%)$
PARADE Attn (ELECTRA) 0.431a $(+3.3\%)$ 0.675a $(+3.5\%)$ 0.653 $(+1.8\%)$ 0.705 $(+2.8\%)$ 0.523a $(+6.4\%)$ 0.581a $(+5.3\%)$
FirstP (DEBERTA) 0.415 0.675 0.629 0.702 0.534 0.596
MaxP (DEBERTA) 0.402 $(-3.2\%)$ 0.679 $(+0.6\%)$ 0.620 $(-1.4\%)$ 0.705 $(+0.4\%)$ 0.535 $(+0.2\%)$ 0.609 $(+2.2\%)$
PARADE Attn (DEBERTA) 0.422a $(+1.6\%)$ 0.685 $(+1.4\%)$ 0.659a $(+4.8\%)$ 0.713 $(+1.4\%)$ 0.549a $(+2.9\%)$ 0.615a $(+3.2\%)$
FirstP (Longformer) 0.404 0.657 0.616 0.654 0.483 0.540
LongP (Longformer) 0.412a $(+1.9\%)$ 0.676a $(+2.9\%)$ 0.628 $(+2.0\%)$ 0.693a $(+6.0\%)$ 0.500a $(+3.6\%)$ 0.568a $(+5.1\%)$
FirstP (Big-Bird) 0.408 0.663 0.620 0.679 0.507 0.560
LongP (Big-Bird) 0.397a $(-2.9\%)$ 0.655 $(-1.1\%)$ 0.618 $(-0.3\%)$ 0.675 $(-0.5\%)$ 0.452a $(-10.9\%)$ 0.477a $(-14.9\%)$
FirstP (JINA) 0.422 0.658 0.618 0.679 0.488 0.532
LongP (JINA) 0.416a $(-1.5\%)$ 0.670a $(+1.8\%)$ 0.632 $(+2.1\%)$ 0.689 $(+1.4\%)$ 0.503a $(+2.9\%)$ 0.558a $(+4.9\%)$
FirstP (MOSAIC) 0.423 0.654 0.607 0.662 0.453 0.538
LongP (MOSAIC) 0.421 $(-0.4\%)$ 0.660 $(+0.9\%)$ 0.630a $(+3.7\%)$ 0.694a $(+4.9\%)$ 0.456 $(+0.6\%)$ 0.570a $(+6.0\%)$
  • In each column we show a relative gain over model’s respective FirstP baseline: The last column shows the average relative gain over FirstP. Best numbers are in bold: Our results are averaged over three seeds. Prior-art seed strategy is unknown.
    Statistical significant differences with respect to this baseline are denoted using the superscript superscript a. $p$-value threshold is 0.01 for an MS MARCO development collection and 0.05 otherwise.
Table 5: Comparison between long-document models and respective FirstP (truncation) baselines for several backbone Transformer models as well as comparison to prior art. Results for MS MARCO, TREC DL, and Robust04.
All MS MARCO models were trained from scratch. Afterward, these models were fine-tuned on Robust04. Note that except for the aggregator Transformers (see D.2 for architectural details) and TinyLLAMA, we use a base, i.e., a 12-layer Transformer Vaswani et al. (2017) models. TinyLLAMA has 22 layers and about 1B parameters.
BERT-base is more practical than a 24-layer BERT-large and performed on par with BERT-large on MS MARCO Documents and Robust04 Hofstätter et al. (2020a); Li et al. (2024). In our own preliminary experiments, we observed that the 24-layer BERT-large model performed much better on the MS MARCO Passage collection, but we were not able to outperform 12-layer BERT-base models on the MS MARCO Documents collection. Note that Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), DEBERTA-base He et al. (2021), JINA Günther et al. (2023), and MOSAIC Portes et al. (2023) all have 12 layers, but a larger embedding matrix compared to BERT-base.
One training epoch consisted in iterating over all queries and sampling one positive and one negative example with a subsequent computation of a pairwise margin loss. We used the mini-batch size one with gradient accumulation over 16 steps. The learning rates are provided in the model configuration files (in the code repository). [^7] https://github.com/searchivarius/long_doc_rank_model_analysis_v2/. We used the AdamW optimizer Loshchilov and Hutter (2017) and a constant learning rate with a 20% linear warm-up Mosbach et al. (2020).
We have learned that—unlike neural retrievers—cross-encoding rankers Nogueira and Cho (2019) are relatively insensitive to learning rates, their schedules, and the choice of loss functions. We were sometimes able to achieve better results using multiple negatives per query and a listwise margin loss (or cross-entropy). However, the gains were small and not consistent compared to a simple pairwise margin loss used in our work (in fact, using a listwise loss function sometimes led to overfitting). Note again that this is different from neural retrievers where training is sometimes difficult without using a listwise loss and/or batch-negatives Karpukhin et al. (2020); Xiong et al. (2021); Qu et al. (2021); Zerveas et al. (2021); Formal et al. (2021).
For MS MARCO, all models except PARADE-Transf-Pretr-LATEIR-L6 and PARADE-Transf-RAND-L2 were trained for one epoch: Further training did not improve (and sometimes degraded) accuracy. However, PARADE-Transf-RAND-L2 and PARADE-Transf-Pretr-LATEIR-L6 required two-to-three epochs to reach the maximum accuracy. For training with the debiased MS MARCO, we used only one epoch. In the case of Robust04, each model was fine-tuned for 100 epochs, but all epochs were short, so the overall training and evaluation time was comparable to that of fine-tuning on MS MARCO for a single epoch.
Except 1B-parameter TinyLLAMA Zhang et al. (2024), each model has about 100M parameters (see Table 11 for details). Despite training and testing 20+ models with three seeds, we estimate to have spent only about 6400 GPU hours for our main experiments. 96% of the time we used NVIDIA A10 (or similarly-powerful) RTX 3090 GPUs and 4% of the time we used NVIDIA A6000.
We believe this is roughly equivalent to pre-training a single 1B-parameter TinyLLAMA model with a next-token prediction objective, which required about 3400 GPU hours using a more powerful NVIDIA A100. This, in turn, this is only a tiny fraction of compute required to train LLAMA2 models (2% compared to a 7B LLAMA2 model). [^8] https://github.com/microsoft/Llama-2-Onnx/blob/main/MODEL-CARD-META-LLAMA-2.md
From our experience, models trained on MS MARCO v2 performed worse on TREC 2021 queries compared to models trained on MS MARCO v1. This may indicate that models implicitly learned to distinguish between original MS MARCO v1 documents and newly added ones (which did not have positive judgements in the training sets). As a result, these models are biased and tend to rank newly added documents poorly even when they are considered to be relevant by NIST assessors. For this reason, we used MS MARCO v2 data only in a zero-shot transfer mode. To this end, ranking models trained on MS MARCO v1 were evaluated using TREC DL 2021 queries.
Retriever / Ranker MS MARCO TREC DL Robust04 Avg. gain
dev (2019-2021) title description Over FirstP
MRR NDCG@10 NDCG@20
BM25 0.274 0.545 0.428 0.402
Retriever (if different from BM25) 0.312 0.629
PARADE Attn (1 chunk) 0.401 0.637 0.476 0.527
PARADE Attn (2 chunks) 0.408a $(+1.8\%)$ 0.653a $(+2.7\%)$ 0.499a $(+4.9\%)$ 0.544a $(+3.3\%)$ $3.2\%$
PARADE Attn (3 chunks) 0.406a $(+1.3\%)$ 0.648a $(+1.7\%)$ 0.505a $(+6.1\%)$ 0.557a $(+5.7\%)$ $3.7\%$
PARADE Attn (4 chunks) 0.412a $(+2.9\%)$ 0.654a $(+2.7\%)$ 0.504a $(+5.9\%)$ 0.558a $(+5.9\%)$ +4.3%
PARADE Attn (5 chunks) 0.409a $(+2.0\%)$ 0.652a $(+2.4\%)$ 0.502a $(+5.6\%)$ 0.556a $(+5.5\%)$ $3.9\%$
PARADE Attn (6 chunks) 0.411a $(+2.4\%)$ 0.653a $(+2.6\%)$ 0.504a $(+5.9\%)$ 0.554a $(+5.2\%)$ $4.0\%$
Table 6: Effectiveness of the PARADE Attention model for different input truncation thresholds. Results for MS MARCO, TREC DL, and Robust04.
A.1.3 Miscellaneous Notes
To enable efficient training and evaluation of the large number of models, both Robust04 and original MS MARCO documents were truncated to have at most 1431 BERT tokens. Thus, for SplitP approaches, queries were padded to 32 BERT tokens and long documents were split into at most three chunks, each containing 477 document tokens. However, for training on debiased MS MARCO, the truncation threshold was much higher: 8109 tokens.
In § B.2.1 (see Table 6) we show that for our top-performing model PARADE Attention Li et al. (2024) using a larger number of chunks only marginally improves outcomes. Depending on a dataset, the highest accuracy is achieved using either three or four chunks.
We evaluated over 20 models, but we had to exclude two LongT5 variants Guo et al. (2022) and RoFormer (with ROPE embeddings) Su et al. (2024) due to poor convergence and/or crashes. Specifically, even after 10 epochs of training LongT5 models were $\approx 10\%$ less accurate than BERT-base FirstP trained for one epoch. Given the uncertainty regarding the possible convergence of models as well as the need to train these for three epochs, we had to abandon this experiment as overly expensive. RoFormer models were failing due to CUDA errors when the context length exceeded 512: We were not able to resolve this issue.
For bias-mitigation experiments, for several reasons, we used only a subset of models. First, we had to exclude all LongP models since none of them supported a context longer than 8192 tokens. In contrast, we trained our chunk-and-aggregate models on documents limited to 8109 tokens and then extrapolated their use to rank documents up to 32768 tokens long.
Second, we chose representative models with vastly different generalization properties. MaxP and PARADE Attention models performed well on MS MARCO FarRelevant in the zero-shot setting, but did not benefit much from in-domain fine-tuning. PARADE Transformer MRR dropped from 0.433 to 0.229 in the zero-shot setting, but increased up to 0.432 after in-domain fine-tuning. CEDR-KNRM also benefited a lot from fine-tuning on MS MARCO FarRelevant, but its zero-shot performance was at the level of the random-baseline.
Appendix B Additional Experimental Results
This section presents supplementary experimental results. In particular, we compute additional effectiveness metrics for MS MARCO, TREC DL, and Robust04. MS MARCO and TREC DL results are shown in Table 9. Results for Robust04 datasets are presented in Table 10. Evaluation results for rankers trained on the debiased MS MARCO and tested on short-document collections are shown in Table 3. Furthermore, we provide detailed results for MS MARCO FarRelevant in Table 4.
To justify the use of BERT-base Devlin et al. (2019) as the main model backbone, we carry out an ablation study in § B.1. Moreover, in this section we also compare our results to prior art, with the objective to boost trustworthiness of our experiments.
This is followed by a miscellaneous ablations subsection (§ B.2), where we carry out an efficiency evaluation and assess if truncating input to at most 1431 BERT tokens affected our experimental outcomes.
B.1 Backbone Selection for SplitP Models and Prior Art Comparison
B.1.1 Choice of a Backbone
To understand if using BERT-base as a backbone model for various SplitP (i.e., chunk-and-aggregate) approaches diminished models’ ability to process and understand long contexts, we carried out a focused comparison of several backbone models, including ELECTRA Clark et al. (2020) and DEBERTA He et al. (2021).
To this end, we used two methods: PARADE Li et al. (2024) Attention and MaxP. PARADE Attention model achieved the largest average gain over FirstP in our main experiments (see Table 1), whereas MaxP models were extensively benchmarked in the past Li et al. (2024); Dai and Callan (2019); Zhang et al. (2021).
Although prior work found ELECTRA to be a better backbone model in terms of absolute accuracy Li et al. (2024); Zhang et al. (2021), we found no systematic evaluation of the relationship between a backbone model and gains achievable over FirstP.
Results in Tables 1 and 5 confirm the overall superiority of both ELECTRA and DEBERTA over BERT-base. In that, DEBERTA models are nearly always more effective compared to ELECTRA models with biggest differences on Robust04. However, their relative effectiveness with respect to their respective FirstP baselines does not exceed that of BERT-base.
The same is true for LongP models. Except Longformer, they performed equally or worse compared to FirstP on 8 test sets out of 18. Moreover, all LongP models achieved lower average gains over FirstP (see the last column in Table 1). We conclude that to measure capabilities of chunk-and-aggregate or long-context models to understand and incorporate long documents, it appears to be beneficial to use BERT-base instead of ELECTRA or DEBERTA.
Finally, we would like to note that on standard benchmarks Big-Bird’s Zaheer et al. (2020) FirstP version always outperformed its LongP version (sometimes by as much as 10-15%), which seems to be puzzling. We noticed, however, that for shorter inputs, the model turns off sparse attention and prints the respective warning. Thus, we hypothesize that it is the use of sparse attention that causes this degradation. In contrast, the sparse attention implementation of the Longformer Beltagy et al. (2020) does not exhibit such a degradation (although with Longformer, not all attention is sparse: query-to-document attention is full). Although Big-Bird underperforms on standard benchmarks, after fine-tuning, it still matches the Longformer accuracy on MS MARCO FarRelevant (see Table 4).
B.1.2 Comparison to Prior Art
We also use Table 5 to compare with prior art. We generally reproduce prior art, in particular, experiments by Li et al. 2024, who invented PARADE models. Our ELECTRA-based models achieve higher NDCG@10 on TREC DL 2019-2020 and PARADE Attention models come very close, but they are about 3-5% worse compared to their PARADE Transformer on Robust04. At the same time, our DEBERTA-based PARADE Attention model achieves similar NDCG@20 scores.
Note that one should not expect identical results due to differences in training regimes and candidate generators. In particular, in the case of Robust04, Li et al. 2024 used RM3, which is BM25 with pseudo-relevance feedback Jaleel et al. (2004). On Robust04, it is more effective than BM25 Robertson (2004), which we used as the first-stage retriever.
Another important comparison point is Robust04 results by Zhang et al. 2021 who were able to reproduce original MaxP results by Dai and Callan (2019), who used BERT-base as a backbone. In addition, they experimented with ELECTRA models “pre-finetuned” on MS MARCO. When comparing results for the BERT-base backbone, Zhang et al. (2021) have the maximum relative gain of 6.6% compared to ours 3.3%. However, in absolute terms we got higher NDCG@20 for both FirstP and MaxP. Their MaxP (ELECTRA) has comparable performance to ours on TREC DL 2019-2020, but it is 2-4% better on Robust04. In turn, our MaxP (DEBERTA) is better by 2-6%. Although we do not always match prior art using the same backbone models, we generally match or outperform prior results, which, we believe, boosts the trustworthiness of our experiments.
Figure 3: Efficiency of long-document models vs respective (truncation) FirstP baselines. The figure shows an average relative gain (in %) vs. relative increase in run-time compared to respective FirstP baselines on MS MARCO, TREC DL 2019-2021, and Robust04 (for a representative subset of models). Except LongP RankGPT, LongP models truncate documents to be at most 1431 tokens. There is no truncation for RankGPT.
B.2 Miscellaneous Ablations
B.2.1 Varying the Number of Chunks
To understand if truncating input to at most 1431 BERT tokens biased our experimental outcomes, we carried out an ablation study where one of the top-performing models was trained and evaluated using inputs of varying maximum lengths. To this end, we used PARADE Attention with the number of input chunks varying from one to six. For each setting, the same number of chunks was used during training and evaluation, i.e., we had to carry out six experiments. Similar to our main experiments, we trained each model using three seeds. We carried out this ablation experiment using our MS MARCO and Robust04 datasets.
The results are presented in Table 6. We can see that—depending on the dataset—three or four input chunks are optimal. However, compared to using three chunks, the additional gains over the FirstP baseline are at most 0.6% when averaged over all test sets.
Gao and Callan 2022 carried out a similar ablation using ClueWeb09. Increasing the number of input chunks from three to six lead to only about 2.3% relative improvement in NDCG@20. However, even this modest gain could have been slightly inflated due to model not being trained directly on shorter inputs. Indeed, truncation of an input for a six-chunk model to one chunk is potentially less effective than training and evaluating the model using one-chunk data.
B.2.2 Efficiency Evaluation
We conducted an efficiency evaluation because long-document models, despite their potential to capture more context, often incur substantial computational overhead compared to their FirstP counterparts, raising questions about their practical utility.
[…]
All LongP models are even slower and show less improvement. Given such small benefits at the cost of a substantial slow-down, one could question practicality of such models and suggest using FirstP variants instead.
Appendix C Additional Dataset Details
C.1 Summary Dataset Statistics
# of queries
avg. # of
BERT tokens
avg. # of
pos. judgements
MS MARCO doc. v1
MS MARCO doc. train 352K 7 1
MS MARCO doc. dev 5193 7 1
TREC DL 2019 43 7 153.4
TREC DL 2020 45 7.4 39.3
MS MARCO v2
TREC DL 2021 57 9.8 143.9
Robust04
title 250 3.6 69.6
description 250 18.7 69.6
MS MARCO FarRelevant
train 50K 7.0 1
test 1K 7.0 1
LongEmbed
Needle 800 13.0 1
Passkey 800 9.7 1
MS MARCO pass. v1
TREC DL 2019 43 7 95.4
TREC DL 2020 54 7.2 66.8
BEIR
Natural Questions 3452 9.9 1.2
Touche 49 19 7.9
TREC COVID 50 493.5 16
NFCorpus (NFC) 323 38.2 5
DBP 400 38.2 6.5
SciFact 300 1.1 20.8
SciDocs 1000 4.9 13.1
Table 7: Query Statistics
data set # of documents average # of
BERT tokens
per document
Long-document collections
MS MARCO doc. v1 3.2M 1.4K
MS MARCO doc. v2 12M 2K
Robust04 0.5M 0.6K
MS MARCO FarRelevant 0.53M 1.1K
Needle (LongEmbed) 0.8K variable-length
Passkey (LongEmbed) 0.8K variable-length
Short-document collections
MS MARCO pass. v1 8.8M 75
Short-document collections: BEIR subset
Natural Questions 2.7M 107
Touche 382.5K 382
TREC COVID 171.3K 243
NFCorpus (NFC) 3.6K 355
DBP 4.6M 71
SciFact 5.2K 335
SciDocs 25.7K 234
  • LongEmbed subsets each have 16 subsets of documents whose lengths vary from (approximately) 256 to 32768 tokens.
Table 8: Document Statistics
Summary query and dataset statistics are shown in Tables 7 and 8. In the case of BEIR, we use seven subsets, which include:
  • Natural Questions Kwiatkowski et al. (2019);
  • Touché 2020 Bondarenko et al. (2020);
  • TREC COVID Voorhees et al. (2021);
  • NFCorpus Boteva et al. (2016);
  • DBPedia-Entity-v2 Hasibi et al. (2017);
  • SciFact Wadden et al. (2020);
  • SciDocs Cohan et al. (2020).
Please, note that in the case of MS MARCO FarRelevant, we created about 500K training and 7K testing queries, but to reduce experimentation cost we ended up using only 50K and 1K queries, respectively.
Model MS MARCO TREC DL
dev 2019-2021
MRR NDCG@10 P@10 MAP
BM25 0.274 0.545 0.636 0.282
Retriever 0.312 0.629 0.720 0.321
FirstP (BERT) 0.394 0.632 0.712 0.311
FirstP (Longformer) 0.404 0.643 0.722 0.317
FirstP (ELECTRA) 0.417 0.662 0.734 0.320
FirstP (DEBERTA) 0.415 0.672 0.741 0.327
FirstP (Big-Bird) 0.408 0.656 0.727 0.321
FirstP (JINA) 0.422 0.654 0.728 0.320
FirstP (MOSAIC) 0.423 0.643 0.726 0.316
FirstP (TinyLLAMA) 0.395 0.615 0.692 0.301
FirstP (E5) 0.380 0.641 0.722 0.317
FirstP RankGPT (OpenAI) 0.708 0.790 0.352
FirstP RankGPT (Anthropic) 0.703 0.776 0.347
AvgP 0.389 $(-1.3\%)$ 0.642 $(+1.5\%)$ 0.717 $(+0.7\%)$ 0.317a $(+2.0\%)$
MaxP 0.392 $(-0.4\%)$ 0.644a $(+1.9\%)$ 0.723 $(+1.5\%)$ 0.322a $(+3.7\%)$
MaxP (ELECTRA) 0.414 $(-0.6\%)$ 0.659 $(-0.5\%)$ 0.745 $(+1.5\%)$ 0.326 $(+2.1\%)$
MaxP (DEBERTA) 0.402a $(-3.2\%)$ 0.671 $(-0.1\%)$ 0.746 $(+0.7\%)$ 0.335a $(+2.5\%)$
SumP 0.390 $(-1.0\%)$ 0.639 $(+1.0\%)$ 0.715 $(+0.4\%)$ 0.319a $(+2.6\%)$
CEDR-DRMM 0.385a $(-2.3\%)$ 0.629 $(-0.5\%)$ 0.708 $(-0.5\%)$ 0.313 $(+0.6\%)$
CEDR-KNRM 0.379a $(-3.8\%)$ 0.630 $(-0.3\%)$ 0.711 $(-0.1\%)$ 0.313 $(+0.8\%)$
CEDR-PACRR 0.395 $(+0.3\%)$ 0.643a $(+1.6\%)$ 0.719 $(+0.9\%)$ 0.320a $(+2.9\%)$
Neural Model1 0.398 $(+0.9\%)$ 0.650a $(+2.8\%)$ 0.723a $(+1.5\%)$ 0.323a $(+3.9\%)$
PARADE Attn 0.416a $(+5.5\%)$ 0.652a $(+3.1\%)$ 0.728a $(+2.2\%)$ 0.324a $(+4.2\%)$
PARADE Attn (ELECTRA) 0.431a $(+3.3\%)$ 0.680a $(+2.7\%)$ 0.763a $(+3.9\%)$ 0.335a $(+4.9\%)$
PARADE Attn (DEBERTA) 0.422a $(+1.6\%)$ 0.688a $(+2.4\%)$ 0.763a $(+3.0\%)$ 0.339a $(+3.9\%)$
PARADE Avg 0.392 $(-0.6\%)$ 0.646a $(+2.1\%)$ 0.715 $(+0.4\%)$ 0.317a $(+2.1\%)$
PARADE Max 0.405a $(+2.7\%)$ 0.655a $(+3.5\%)$ 0.733a $(+2.9\%)$ 0.324a $(+4.1\%)$
PARADE Transf-RAND-L2 0.419a $(+6.3\%)$ 0.655a $(+3.6\%)$ 0.734a $(+3.1\%)$ 0.326a $(+5.0\%)$
PARADE Transf-RAND-L2 (ELECTRA) 0.433a $(+3.9\%)$ 0.670 $(+1.2\%)$ 0.747 $(+1.8\%)$ 0.327 $(+2.2\%)$
PARADE Transf-PRETR-L6 0.402a $(+1.9\%)$ 0.643 $(+1.6\%)$ 0.717 $(+0.8\%)$ 0.322a $(+3.6\%)$
PARADE Transf-PRETR-LATEIR-L6 0.398 $(+1.1\%)$ 0.626 $(-0.9\%)$ 0.707 $(-0.7\%)$ 0.307 $(-1.1\%)$
LongP (Longformer) 0.412a $(+1.9\%)$ 0.668a $(+3.9\%)$ 0.752a $(+4.1\%)$ 0.333a $(+5.1\%)$
LongP (Big-Bird) 0.397a $(-2.9\%)$ 0.651 $(-0.7\%)$ 0.726 $(-0.2\%)$ 0.322 $(+0.3\%)$
LongP (JINA) 0.416a $(-1.5\%)$ 0.665a $(+1.7\%)$ 0.742a $(+2.0\%)$ 0.328a $(+2.4\%)$
LongP (MOSAIC) 0.421 $(-0.4\%)$ 0.664a $(+3.3\%)$ 0.740a $(+1.9\%)$ 0.327a $(+3.7\%)$
LongP (TinyLLAMA) 0.402a $(+1.7\%)$ 0.608 $(-1.1\%)$ 0.692 $(+0.0\%)$ 0.306 $(+1.6\%)$
LongP (E5) 0.353a $(-7.1\%)$ 0.649 $(+1.3\%)$ 0.724 $(+0.3\%)$ 0.323 $(+1.8\%)$
LongP RankGPT (OpenAI) 0.706 $(-0.3\%)$ 0.783 $(-1.0\%)$ 0.350 $(-0.7\%)$
LongP RankGPT (Anthropic) 0.707 $(+0.5\%)$ 0.780 $(+0.5\%)$ 0.348 $(+0.4\%)$
  • In each column we show a relative gain with respect model’s respective FirstP baseline: The last column shows the average relative gain of FirstP. Best numbers are in bold: Results are averaged over three seeds. Unless specified explicitly, the backbone is BERT-base.
    Statistical significant differences with respect to this baseline are denoted using the superscript superscript a. $p$-value threshold is 0.01 for an MS MARCO development collection and 0.05 otherwise.
    E5-models were used only in the zero-shot model, i.e., without fine-tuning.
Table 9: Comparison between long-document models and respective FirstP (truncation) baselines. Results for MS MARCO and TREC DL.
Model NDCG@20 P@20 MAP NDCG@20 P@20 MAP
Retriever (BM25) 0.428 0.365 0.255 0.402 0.334 0.240
FirstP (BERT) 0.475 0.405 0.277 0.527 0.447 0.303
FirstP (Longformer) 0.483 0.413 0.277 0.540 0.454 0.307
FirstP (ELECTRA) 0.492 0.424 0.294 0.552 0.465 0.320
FirstP (DEBERTA) 0.534 0.459 0.319 0.596 0.503 0.350
FirstP (Big-Bird) 0.507 0.435 0.300 0.560 0.473 0.325
FirstP (JINA) 0.488 0.421 0.287 0.532 0.450 0.305
FirstP (MOSAIC) 0.453 0.390 0.266 0.538 0.455 0.310
FirstP (TinyLLAMA) 0.431 0.370 0.246 0.473 0.398 0.262
FirstP (E5-4K) 0.438 0.371 0.247 0.429 0.355 0.234
FirstP RankGPT (OpenAI) 0.562 0.456 0.280
FirstP RankGPT (Anthropic) 0.541 0.446 0.268
AvgP 0.478 $(+0.5\%)$ 0.411 $(+1.6\%)$ 0.292a $(+5.4\%)$ 0.531 $(+0.9\%)$ 0.451 $(+1.0\%)$ 0.324a $(+6.7\%)$
MaxP 0.488a $(+2.6\%)$ 0.425a $(+5.1\%)$ 0.306a $(+10.6\%)$ 0.544a $(+3.3\%)$ 0.467a $(+4.5\%)$ 0.338a $(+11.5\%)$
MaxP (ELECTRA) 0.502 $(+2.0\%)$ 0.441a $(+3.9\%)$ 0.319a $(+8.3\%)$ 0.563 $(+2.1\%)$ 0.483a $(+4.0\%)$ 0.350a $(+9.3\%)$
MaxP (DEBERTA) 0.535 $(+0.2\%)$ 0.464 $(+1.2\%)$ 0.340a $(+6.7\%)$ 0.609a $(+2.2\%)$ 0.519a $(+3.2\%)$ 0.378a $(+7.9\%)$
SumP 0.486 $(+2.2\%)$ 0.418a $(+3.4\%)$ 0.305a $(+10.2\%)$ 0.538 $(+2.1\%)$ 0.461a $(+3.1\%)$ 0.337a $(+11.1\%)$
CEDR-DRMM 0.466 $(-2.0\%)$ 0.403 $(-0.4\%)$ 0.287a $(+3.8\%)$ 0.533 $(+1.3\%)$ 0.458a $(+2.5\%)$ 0.326a $(+7.6\%)$
CEDR-KNRM 0.483 $(+1.7\%)$ 0.413 $(+1.9\%)$ 0.291a $(+5.1\%)$ 0.535 $(+1.7\%)$ 0.456 $(+2.0\%)$ 0.324a $(+6.8\%)$
CEDR-PACRR 0.496a $(+4.3\%)$ 0.426a $(+5.3\%)$ 0.307a $(+11.0\%)$ 0.549a $(+4.2\%)$ 0.466a $(+4.4\%)$ 0.337a $(+11.2\%)$
Neural Model1 0.484 $(+1.8\%)$ 0.417a $(+3.1\%)$ 0.298a $(+7.7\%)$ 0.537 $(+1.9\%)$ 0.459a $(+2.6\%)$ 0.330a $(+8.8\%)$
PARADE Attn 0.503a $(+5.7\%)$ 0.433a $(+6.9\%)$ 0.311a $(+12.4\%)$ 0.556a $(+5.6\%)$ 0.476a $(+6.5\%)$ 0.344a $(+13.3\%)$
PARADE Attn (ELECTRA) 0.523a $(+6.4\%)$ 0.456a $(+7.4\%)$ 0.329a $(+11.7\%)$ 0.581a $(+5.3\%)$ 0.495a $(+6.5\%)$ 0.358a $(+11.9\%)$
PARADE Attn (DEBERTA) 0.549a $(+2.9\%)$ 0.475a $(+3.6\%)$ 0.346a $(+8.7\%)$ 0.615a $(+3.2\%)$ 0.522a $(+3.8\%)$ 0.383a $(+9.4\%)$
PARADE Avg 0.483 $(+1.5\%)$ 0.412 $(+1.8\%)$ 0.291a $(+5.2\%)$ 0.534 $(+1.5\%)$ 0.457 $(+2.4\%)$ 0.318a $(+4.7\%)$
PARADE Max 0.489a $(+2.8\%)$ 0.420a $(+3.8\%)$ 0.306a $(+10.8\%)$ 0.548a $(+4.0\%)$ 0.470a $(+5.3\%)$ 0.337a $(+11.0\%)$
PARADE Transf-RAND-L2 0.488a $(+2.8\%)$ 0.423a $(+4.6\%)$ 0.303a $(+9.7\%)$ 0.548a $(+4.1\%)$ 0.469a $(+5.0\%)$ 0.338a $(+11.6\%)$
PAR. Transf-RAND-L2 (ELECTRA) 0.523a $(+6.3\%)$ 0.454a $(+6.9\%)$ 0.330a $(+12.2\%)$ 0.574a $(+3.9\%)$ 0.488a $(+5.0\%)$ 0.354a $(+10.6\%)$
PARADE Transf-PRETR-L6 0.494a $(+4.0\%)$ 0.426a $(+5.3\%)$ 0.308a $(+11.5\%)$ 0.554a $(+5.1\%)$ 0.474a $(+6.1\%)$ 0.346a $(+14.1\%)$
PAR. Transf-PRETR-LATEIR-L6 0.450a $(-5.2\%)$ 0.389a $(-3.9\%)$ 0.277 $(+0.3\%)$ 0.501a $(-4.9\%)$ 0.423a $(-5.3\%)$ 0.302 $(-0.5\%)$
LongP (Longformer) 0.500a $(+3.6\%)$ 0.435a $(+5.3\%)$ 0.309a $(+11.5\%)$ 0.568a $(+5.1\%)$ 0.482a $(+6.1\%)$ 0.347a $(+12.9\%)$
LongP (Big-Bird) 0.452a $(-10.9\%)$ 0.389a $(-10.7\%)$ 0.274a $(-8.8\%)$ 0.477a $(-14.9\%)$ 0.400a $(-15.5\%)$ 0.279a $(-14.2\%)$
LongP (JINA) 0.503a $(+2.9\%)$ 0.434a $(+3.1\%)$ 0.309a $(+7.5\%)$ 0.558a $(+4.9\%)$ 0.473a $(+5.2\%)$ 0.335a $(+9.7\%)$
LongP (MOSAIC) 0.456 $(+0.6\%)$ 0.393 $(+0.8\%)$ 0.280a $(+5.3\%)$ 0.570a $(+6.0\%)$ 0.484a $(+6.3\%)$ 0.350a $(+13.0\%)$
LongP (TinyLLAMA) 0.452a $(+4.8\%)$ 0.396a $(+6.9\%)$ 0.267a $(+8.7\%)$ 0.505a $(+6.7\%)$ 0.428a $(+7.6\%)$ 0.297a $(+13.3\%)$
LongP (E5-4K) 0.439 $(+0.1\%)$ 0.375 $(+1.0\%)$ 0.250 $(+1.3\%)$ 0.434 $(+1.1\%)$ 0.360 $(+1.6\%)$ 0.241a $(+2.9\%)$
LongP RankGPT (OpenAI) 0.562 $(+0.0\%)$ 0.456 $(-0.0\%)$ 0.281 $(+0.3\%)$
LongP RankGPT (Anthropic) 0.538 $(-0.6\%)$ 0.445 $(-0.2\%)$ 0.268 $(-0.0\%)$
  • In each column we show a relative gain with respect model’s respective FirstP baseline: The last column shows the average relative gain of FirstP. Best numbers are in bold: Results are averaged over three seeds. Unless specified explicitly, the backbone is BERT-base.
    Statistical significant differences with respect to this baseline are denoted using the superscript superscript a. $p$-value threshold is 0.05.
    E5-models were used only in the zero-shot model, i.e., without fine-tuning.
Table 10: Comparison between long-document models and respective FirstP (truncation) baselines. Results for Robust04.
C.2 MS MARCO FarRelevant Creation Algorithm
The algorithm iterates over the set of relevant passages in the original MS MARCO Passages collection. Each iteration is as follows:
  1. 1.
    Assume that $C_{t}$ is the number of tokens in the selected relevant passage. Select randomly a document length $D_{max}$ between $512+C_{t}$ and 1431;
  2. 2.
    Using rejection sampling, obtain $K_{1}$ non-relevant passages such that their total length is $\geq 512$, but the length of $K_{1}-1$ first passages is <512<512.
  3. 3.
    Using the same approach, sample another $K_{2}$ non-relevant passages such that the total length of $K_{1}+K_{2}-1$ non-relevant passages is at most $D_{max}-C_{t}$, but the total length of $K_{1}+K_{2}$ passages is $\geq D_{max}-C_{t}$.
  4. 4.
    Randomly mix these $K_{2}$ non-relevant passages with a single relevant passage.
  5. 5.
    Finally, append the resulting string to the concatenation of the first $K_{1}$ non-relevant passages.
Initially, we planned to limit documents to have at most 1431 tokens. However, step 3 in the above algorithm has a bug: It samples one extra non-relevant passage. As a result, only about 5% of generated documents have more than 1431 tokens and less than 1% have more than 1500.
{textenv}
Aaron Swartz created a scraped feed of the essays page. November 2021(This essay is derived from a talk at the Cambridge Union. )When I was a kid, I’d have said there wasn’t. My father told me so. Some people like some things, and other people like other things, and who’s to say who’s right?It seemed so obvious that there was no such thing as good taste that it was only through indirect evidence that I realized my father was wrong. And that’s what I’m going to give you here: a proof by reductio ad absurdum. If we start from the premise that there’s no such thing as good taste, we end up with conclusions that are obviously false, and therefore the premise must be wrong. We’d better start by saying what good taste is. There’s a narrow sense in which it refers to aesthetic judgements and a broader one in which it refers to preferences of any kind. The strongest proof would be to show that taste exists in the narrowest sense, so I’m going to talk about taste in art. You have better taste than me if the art you like is better than the art I like. If there’s no such thing as good taste, then there’s no such thing as good art. Because if there is such a thing as good art, it’s easy to tell which of two people has better taste. Show them a lot of works by artists they’ve never seen before and ask them to choose the best, and whoever chooses the better art has better taste. So if you want to discard the concept of good taste, you also have to discard the concept of good art. And that means you have to discard the possibility of people being good at making it. Which means there’s no way for artists to be good at their jobs. And not just visual artists, but anyone who is in any sense an artist. You can’t have good actors, or novelists, or composers, or dancers either. You can have popular novelists, but not good ones. We don’t realize how far we’d have to go if we discarded the concept of good taste, because we don’t even debate the most obvious cases. But it doesn’t just mean we can’t say which of two famous painters is better. It means we can’t say that any painter is better than a randomly chosen eight year old. That was how I realized my father was wrong. I started studying painting. And it was just like other kinds of work I’d done: you could do it well, or badly, and if you tried hard, you could get better at it. And it was obvious that Leonardo and Bellini were much better at it than me. That gap between us was not imaginary. They were so good. And if they could be good, then art could be good, and there was such a thing as good taste after all. Now that I’ve explained how to show there is such a thing as good taste, I should also explain why people think there isn’t. There are two reasons. One is that there’s always so much disagreement about taste. Most people’s response to art is a tangle of unexamined impulses. Is the artist famous? Is the subject attractive? Is this the sort of art they’re supposed to like? Is it hanging in a famous museum, or reproduced in a big, expensive book? In practice most people’s response to art is dominated by such extraneous factors.
The Terracotta Army is a collection of terracotta sculptures depicting the armies of Qin Shi Huang, the first Emperor of China.
And the people who do claim to have good taste are so often mistaken. The paintings admired by the so-called experts in one generation are often so different from those admired a few generations later. It’s easy to conclude there’s nothing real there at all. It’s only when you isolate this force, for example by trying to paint and comparing your work to Bellini’s, that you can see that it does in fact exist. The other reason people doubt that art can be good is that there doesn’t seem to be any room in the art for this goodness. The argument goes like this. Imagine several people looking at a work of art and judging how good it is. If being good art really is a property of objects, it should be in the object somehow. But it doesn’t seem to be; it seems to be something happening in the heads of each of the observers. And if they disagree, how do
Figure 4: A sample relevant document for the Needle collection. The query/question is: “What is the Terracotta Army?”. The answer-bearing sentence is marked by bold font.
{textenv}
The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green.
Jimmy Moses’s pass key is 39566. Remember it. 39566 is the pass key for Jimmy Moses.
The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again. The grass
Figure 5: A sample relevant document for the Passkey collection. The query/question is: “what is the passkey for Jimmy Moses?”. The answer-bearing sentence is marked by bold font.
{textenv}
Andhra Pradesh Airports make an easy access for tourists visiting the state. This huge state has many airports, which serves the needs of both tourists and residents commuting to different parts in its large expanse. However, Hyderabad Airport is the major as well as the only international airport of Andhra Pradesh. Hyderabad, a major IT hub of India, boasts of the sixth busiest airport in India. Keeping the rush of passengers in mind, the Government is planning to establish another airport in Hyderabad.
In contrast, traditional English Longbow shooters step into the bow, exerting force with both the bow arm and the string hand arm simultaneously, especially when using bows having draw weights from 100 lbs to over 175 lbs. Heavily stacked traditional bows (recurves, long bows, and the like) are released immediately upon reaching full draw at maximum weight, whereas compound bows reach their maximum weight around the last inch and a half, dropping holding weight significantly at full draw. The Oreo Biscuit was first developed and produced by the National Biscuit Company (today known as Nabisco) in 1912 at its Chelsea, Manhattan factory in the current-day Chelsea Market complex, located on Ninth Avenue between 15th and 16th Streets. Today, this same block of Ninth Avenue is known as Oreo Way..
It’s possible to take a day trip to the Bahamas by ferry. In some cases, less than it would cost to fly. The high-speed Balearia Bahamas Express travels from Fort Lauderdale to the city of Freeport on Grand Bahama, one of many Bahamian islands with British roots.It’s roughly the distance from Philadelphia to New York.Prepare for a long day.n some cases, less than it would cost to fly. The high-speed Balearia Bahamas Express travels from Fort Lauderdale to the city of Freeport on Grand Bahama, one of many Bahamian islands with British roots. It’s roughly the distance from Philadelphia to New York.
I was 17 when I took Bactrim for a UTI, I was a month away from turning 18 and was given this because another antibiotic would have more side effects I was told. I took it for 5 days and became horribly sick from a nasty cold and was told to stop Bactrim and to take Z Pac instead for 5 days.
Cases When Medicare Does NOT Automatically Start for You Medicare will NOT automatically start when you turn 65 if you’re not receiving Social Security Benefits or Railroad Retirement Benefits for at least 4 months prior to your 65th birthday. You wanted to know how you can feel on your belly that you’re pregnant. And actually this a pretty hard thing to do.There are better ways to find out if you’re pregnant or not. For example, if you ever miss a period, the best thing to do is to take a home pregnancy test, because that’s the first sign of pregnancy.And if it’s positive, …here are better ways to find out if you’re pregnant or not. For example, if you ever miss a period, the best thing to do is to take a home pregnancy test, because that’s the first sign of pregnancy.
1 Whisk light soy sauce, dark soy sauce, red wine vinegar, chili oil, ginger, sugar, garlic, and green onion together in a bowl; pour into a sealable container, seal, and refrigerate 1 hour. See how to make a simple sweet-and-sour peach sauce. See how easy and delicious it is to make horseradish sauce from scratch.
Here you’ll find a number of Kentucky facts including the state history at a glance; Kentucky state facts such as the location of the state capital, city populations, geography and natural resources; Information on Kentucky’ government, symbols and traditions; and even a list of famous Kentuckians.
Confidence votes 193. Replacing a car window usually cost around \$300 give or take based upon where you live, what options your car glass needs and what type of car you drive. Additionally, you can save money if you can repair your car glass instead of completely replacing it. However, if the window is completely shattered this will not be an option.
Flying time from Chicago, IL to Cairo, Egypt. The total flight duration from Chicago, IL to Cairo, Egypt is 12 hours, 47 minutes. This assumes an average flight speed for a commercial airliner of 500 mph, which is equivalent to 805 km/h or 434 knots. It also adds an extra 30 minutes for take-off and landing. Your exact time may vary depending on wind speeds.
Figure 6: A sample relevant document for the MS MARCO FarRelevant collection. The query/question is: “how long is the flight from chicago to cairo”. The answer-bearing passage is marked by bold font.
C.3 Comparison of FarRelevant and Synthetic Data from LongEmbed
Our synthetic data consists of two subsets (Needle and Passkey) from the LongEmbed collection Zhu et al. (2024) and our newly created MS MARCO FarRelevant dataset. All these datasets can be seen a variant of a needle-in-the-haystack benchmark Saad-Falcon et al. (2024); Zhu et al. (2024). They share a common limitation: the resulting documents are constructed by combining pieces of text in a purely mechanical fashion and do not represent natural documents. In this section, we describe Needle and Passkey in more detail and compare them to our proposed collection MS MARCO FarRelevant, which—we believe—offers several practical advantages despite also being synthetic.
The Needle subset was created by taking a single document (a Paul Graham essay on taste [^9] https://www.paulgraham.com/goodtaste.html), truncating it to generate 16 text pieces of varying lengths (from 256 to 32,768 tokens), and inserting a single answer-bearing sentence at a random location. An example query-document pair can be found in Figure 4. While this design ensures precise control over document length, it introduces several problems:
  1. 1.
    Extremely low document diversity: All examples are truncated variants of the same original text, which severely limits the variability in document style and content.
  2. 2.
    Artificial signal separation: The inserted answer sentence differs substantially from the background text, making it easy for models to identify it.
The Passkey subset is similar in structure and also uses length-bucketed documents, but instead of a single sentence, it inserts a three-sentence passkey definition into a synthetic background context (see Figure 5). The background is even less natural than Needle’s subset background, being composed of unrelated or nonsensical declarative statements (e.g., “The sky is blue. The sun is yellow. Here we go. There and back again.”). This leads to even greater distributional mismatch between signal and context.
Critically, both Needle and Passkey were designed primarily to test answer extraction or retrieval of small, highly localized answer-bearing spans. They are not well suited to studying retrieval or ranking of entire passages or documents, especially when relevance is more distributed or contextual.
Our MS MARCO FarRelevant (see Figure 6 for an example) is designed to be textually similar to MS MARCO Documents but with different positional biases for relevant passages. We believe it offers a more robust testbed for long-context document ranking. While it is also synthetic in construction, it avoids many of the pitfalls noted above. Each document is created by concatenating multiple passages, which are typically meaningful and complete, only one of which is relevant to the query. The remaining passages serve as distractors but are independently coherent. Although these documents do not exist in the wild, they are much more diverse in content and style than those in Needle or Passkey despite being typically much shorter. Furthermore, each individual passage is semantically complete and belongs to a real corpus, namely, MS MARCO Passages.
Crucially, one may argue that it is natural for the models to fail on randomly concatenated passages because they do not represent a complete “natural” document (this criticism applies to the synthetic subsets of LongEmbed as well). Although this is possible for LongP methods—where longer document lengths are “natively” supported by models—nearly all models in our study are SplitP models, which chunk documents and process each chunk separately. Hence, these models do not “care” whether the full document is coherent or not. Furthermore, the strong performance of all the ranking models after fine-tuning on MS MARCO FarRelevant demonstrates that models do not inherently fail on unnatural documents.
C.4 Positional Bias Identification
(a) MS MARCO train
(b) MS MARCO dev
(c) TREC DL 2019-2021 (combined)
(d) Gov2 (1MQ)
(e) ClueWeb12 (WebTrack)
(f) Robust04
Figure 7: Illustration of positional relevance bias for four long-document collections and six query sets. We show a distribution of first relevant passage positions (red bars) vs. relevant document lengths (blue bars). Positions and lengths are measured in the number of subword tokens (BERT-base tokenizer). Best viewed in color.
(a) Natural Questions (NQ)
(b) Touche
(c) TREC COVID
(d) NFCorpus (NFC)
(e) SciFact
(f) SciDocs
Figure 8: Illustration of positional relevance bias for six BEIR collections, all of which are typically categorized as short-document datasets. We show a distribution of first relevant passage positions (red bars) vs. relevant document lengths (blue bars). Positions and lengths are measured in the number of subword tokens (BERT-base tokenizer). Best viewed in color.
To assess positional bias, we used a combination of approximate string matching and LLM-based judging, which was recently shown to highly correlate with human judgments Upadhyay et al. (2024); Arabzadeh and Clarke (2025). The resulting distributions can be found in Figures 7 and 8.
Approximate string matching was used for MS MARCO training and development (dev) sets, both of which have sparse labels. Although initially MS MARCO Passages were exact substrings of MS MARCO Documents, document and passage texts were collected at different times. This led to some content divergence Craswell et al. (2021a), which made exact mapping impossible in most cases. There were prior attempts to recover initial positions, but only with a limited success (see a discussion below).
Our approximate string matching combines approximate substring matching with longest-substring matching and incorporates efficiency heuristics to identify initial candidate sets. Candidate sets were constructed using two approaches: selecting all relevant passages and documents (for a given query) and retrieving top-5 documents using relevant passages as queries. To assess reliability, we manually inspected a subset of the matched passages and found the procedure to be sufficiently accurate. We then applied this approach to two sets of queries:
  • A set of all 5193 queries from the devevelopment set;
  • A random and uniform sample of 5000 training queries.
In both cases, we were able to find matches for about 85% of the queries.
For queries sets with “dense” relevance judgments—produced by TREC NIST assessors—we used an LLM judge. This included TREC 2019, 2020, 2021 TREC DL queries Craswell et al. (2021b), Robust04 Clarke et al. (2004), Gov2 with 2007, 2008 Million Query Track queries Allan et al. (2008), and ClueWeb12 with 2012 and 2013 Web Track queries Collins-Thompson et al. (2013b). We also used an LLM judge for BEIR datasets Thakur et al. (2021).
In contrast, Hofstätter et al. (2020b) employed crowd-workers to identify the distribution of relevance chunks. They found similar evidence of the relevance position bias, but their study was limited only to the TREC DL 2019 query set. For sparsely-judged training queries, Hofstätter et al. (2020b) used matching on answer words—rather than passage text itself—and were able to match only 32% of the passages. In particular, Coelho et al. (2024) used exact matching and found only about one thousand matching passages. Both of these attempts match only a small-to-modest fraction of passages while being subject to biases. Finally, MS MARCO v2 authors Craswell et al. (2021b) provided explicit mappings of relevant passages inside documents, but how these mappings were obtained was not disclosed.
To avoid potential positional biases in LLM-judging, we divided each document into non-overlapping chunks and judged each chunk separately. Chunking preserved sentence boundary while ensuring each chunk size was close to 256 tokens in length for long-document collections and 128 tokens for BEIR collections Thakur et al. (2021).
For efficiency reasons, we only considered at most 36 chunks (about 9K tokens) and 500-2000 positive query-document pairs per query set. A chunk was considered to be relevant if it received a positive grade from an LLM-judge. The LLM-judge was a GPT4-mini-based UMBRELA OpenAI (2023); Upadhyay et al. (2024).
For most collections, there was only a small fraction of documents where the LLM-judge found no relevant query-document pairs. Among long-document datasets, one exception was the Gov2 collection with 2007, 2008 Million Query Track queries where this happened in about 30% of the cases.
Despite this uncertainty, Gov2 still had a substantial positional bias with about 57% of the cases where the first chunk was deemed to be relevant (see Fig. 7). Among BEIR datasets, five datasets also had a prominent positional bias (see Fig. 8). However, we are less confident about the NFCorpus, where we found no relevant chunks for nearly 50% of the query-document pairs.
Model family # of
params.
PARADE Transformer 132-148M
Longformer 149M
BigBird 127M
JINA 137M
MOSAIC 137M
DEBERTA-based models 184M
TinyLLAMA-based models 1034M
Other BERT- and ELECTRA-based models $\approx$110 M
Table 11: Number of model parameters (we do not include models behind cloud APIs)
Appendix D Ranking with Cross-Encoding Long-Document Models
In this section, we describe long-document cross-encoding models in more details. With an exception of TinyLLAMA Zhang et al. (2024) all models employ relatively small bidirectional encoder-only Transformers Vaswani et al. (2017) with 100-200M parameters in total (see Table 11). TinyLLAMA is a so-called LLM-ranker: a “causal” decoder-only Transformer that has about 1B parameters. Moreover, we focus exclusively on pure Transformer architectures, leaving hybrid approaches such as RankMamba Xu (2024) for future work.
We assume that an input text is split into small chunks of texts called tokens. Although tokens can be complete English words, Transformer models usually split text into sub-word units Wu et al. (2016).
The length of a document dd—denoted as |d||d|—is measured in the number of tokens. Because neural networks cannot operate directly on text, a sequence of tokens $t_{1}t_{2}\ldots t_{n}$ is first converted to a sequences of $d$-dimensional embedding vectors $w_{1}w_{2}\ldots w_{n}$ by an embedding network. These embeddings are context-independent, i.e., each token is always mapped to the same vector Collobert et al. (2011); Mikolov et al. (2013).
For a detailed description of Transformer models, please see the annotated Transformer guide Rush (2018) as well as the recent survey by Lin et al. Lin (2019), which focuses on the use of BERT-style cross-encoding models for ranking and retrieval. For this paper, it is necessary to know only the following basic facts:
  • BERT is an encoder-only model, which converts a sequence of tokens $t_{1}t_{2}\ldots t_{n}$ to a sequence of $d$-dimensional vectors $w_{1}w_{2}\ldots w_{n}$. These vectors—which are token representations from the last model layer—are commonly referred to as contextualized token embeddings Peters et al. (2018);
  • BERT operates on word pieces Wu et al. (2016) rather than on complete words;
  • The vocabulary includes two special tokens: [CLS] (an aggregator) and [SEP] (a separator);
  • Using a pooled representation of token vectors $w_{1}w_{2}\ldots w_{n}$, a linear layer is used to produce a ranking score. A nearly universal pooling approach in cross-encoding rankers is to use the vector of the [CLS] token, i.e., $w_{1}$. However, we learned that some models (e.g., JINA Günther et al. (2023)) converge well only with mean pooling, i.e., they use ${1\over n}\sum_{i=1}^{n}w_{i}$.
A “vanilla” BERT ranker—dubbed as monoBERT by Lin (2019)—uses a single fully-connect layer $F$ as a prediction head, which converts the last-layer representation of the [CLS] token (i.e., a contextualized embedding of [CLS]) into a scalar Nogueira and Cho (2019). It makes a prediction based on the following sequence of tokens: [CLS] $q$ [SEP] $d$ [SEP], where $q$ is a query and $d$ is a document.
An alternative approach is to aggregate contextualized embeddings of regular tokens using a shallow neural network MacAvaney et al. (2019); Boytsov and Kolter (2021); Khattab and Zaharia (2020), optionally combining these embeddings with the contextualized embedding of [CLS].
This was first proposed by MacAvaney et al. (2019) who also found that incorporating [CLS] improved performance. However, Boytsov and Kolter (2021) proposed a shallow aggregating network that did not use the output of the [CLS] token and achieved the same accuracy on MS MARCO datasets, thus questioning the necessity of using [CLS] for aggregation.
Replacing the standard BERT model in the vanilla BERT ranker with a BERT variant that “natively” supports longer documents is conceptually the simplest way to handle long documents. We collectively refer to these models as LongP models. For many older BERT models with limited context size, long documents and queries still need to be split or truncated so that the total number of tokens does not exceed 512. In the FirstP mode, we process only the first chunk and ignore the truncated text. In the SplitP mode, each chunk is processed separately and the results are aggregated. In the remaining of this section, we discuss LongP and SplitP approaches in detail. We conclude the discussion with a description of several miscellaneous models, most of which we were unable to implement due to incomplete information or missing code.
D.1 LongP models
In our work, we benchmark both sparse-attention and full-attention models. Sparse attention LongP models include two popular variants: Longformer Beltagy et al. (2020) and Big-Bird Zaheer et al. (2020). In this setting, we use the same approach to score documents as with the vanilla BERT ranker, namely, concatenating queries with documents and making a prediction based on the contextualized embedding of the [CLS] token Nogueira and Cho (2019). Both Big-Bird and Longformer use a combination of the local, “scattered” (our terminology), and global attention. The local attention utilizes a sliding window of a constant length where each token attends to each other token within this window. In the case of the global attention, certain tokens can attend to all other tokens and vice-versa, In Big-Bird, only special tokens such as [CLS] can attend globally. In Longformer, the user have to select such tokens explicitly. Following Beltagy et al. Beltagy et al. (2020), who applied this technique to question-answering, we designate global attention only for query tokens. Unlike the global attention, the scattered attention is limited to restricted sub-sets of tokens, but these subsets do not necessarily have locality. In Big-Bird, the scattered attention uses randomly selected tokens, whereas Longformer uses a dilated sliding-window attention with layer- and head-specific dilation.
Full-attention models include JINABert Günther et al. (2023), TinyLLAMA Zhang et al. (2024), and MosaicBERT Portes et al. (2023), henceforth, simply JINA, TinyLLAMA and MOSAIC. All these models use a recently proposed FlashAttention Dao et al. (2022) to efficiently process long-contexts as well as special positional embeddings that can generalize to document lengths not seen during training. In particular, JINA and MOSAIC use AliBi Press et al. (2022), while TinyLLAM uses ROPE embeddings Su et al. (2023). JINA and MOSAIC are bidirectional encoder-only Transformer model whereas TinyLLAMA is a unidirectional (sometimes called causal) decoder-only Transformer model Vaswani et al. (2017).
Beyond architectural variations, models also vary in their pretraining strategies. MOSAIC relies primarily on the masked language (MLM) objective without next sentence prediction (NSP). JINA uses this approach as a first step and fine-tuning on retrieval and classification tasks with mean-pooled representations as a second step (following a RoBERTa pretraining strategy Liu et al. (2019)). TinyLLAMA was trained using an autoregressive language modeling objective Zhang et al. (2024). We found that JINA lost an ability to effectively pool on the [CLS] token and we used mean-pooling instead. We also use mean pooling for TinyLLAMA. For MOSAIC we used pooling on [CLS].
D.2 SplitP models
SplitP models differ in partitioning and aggregation approaches. Documents can be split into either disjoint or overlapping chunks. In the first case, documents are split in a greedy fashion so that each document chunk—except possibly the last one—is exactly 512 tokens long (after being concatenated with a (padded) query and three special tokens). Because we pad queries to be 32 tokens long, each document chunk contains at most 512323=477512-32-3=477 tokens.
In the second case, we use a sliding window approach with a window size and stride that are not tied to the maximum length of BERT input. Because our primary focus is accuracy and we aim to understand the limits of long-document models, we exclude from evaluation several SplitP models, which achieve better efficiency-effectiveness trade-offs by pre-selecting certain document parts and feeding only selected parts into a BERT ranker. This includes, but is not limited to, papers by Hofstätter et al. (2021b); Zou et al. (2021).
Greedy partitioning into disjoint chunks
CEDR models MacAvaney et al. (2019) and the Neural Model 1 Boytsov and Kolter (2021) use the first method, which involves:
  • tokenizing the document $d$;
  • greedily splitting a tokenized document $d$ into $m$ disjoint chunks: $d=d_{1}d_{2}\ldots d_{m}$;
  • generating $m$ token sequences [CLS] $q$ [SEP] $d_{i}$ [SEP] by concatenating the query with document chunks;
  • processing each sequence with a BERT model to generate contextualized embeddings for regular tokens as well as for [CLS].
The outcome of this procedure is $m$ [CLS]-vectors $cls_{i}$ and $n$ contextualized vectors $w_{1}w_{2}\ldots w_{n}$ (one for each document token $t_{i}$) that are aggregated in a model-specific way.
MacAvaney et al. (2019) use contextualized embeddings as a direct replacement of context-free embeddings in the following neural architectures: KNRM Xiong et al. (2017), PACRR Hui et al. (2018), and DRMM Guo et al. (2016). To boost performance, they incorporate [CLS] -vectors in a model-specific way. We call the respective models as CEDR-KNRM, CEDR-PACRR, and CEDR-DRMM.
They also proposed an extension of the vanilla BERT ranker that makes a prediction using the average [CLS] token: $\frac{1}{m}\sum_{i=1}^{m}cls_{i}$ by passing it through a linear projection layer. We call this method AvgP.
The Neural Model 1 Boytsov and Kolter (2021) uses the same greedy partitioning approach as CEDR, but a different aggregator network, which does not use the embeddings of the [CLS] token. This network is a neural parametrization of the classic Model 1 Berger and Lafferty (1999); Brown et al. (1993).
Sliding window approach
The BERT MaxP/SumP Dai and Callan (2019) and PARADE Li et al. (2024) models use a sliding window approach. Assume $w$ is the size of the window and $s$ is the stride. Then the processing can be summarized as follows:
  • tokenizing, the document $d$ into sub-words $t_{1}t_{2}\ldots t_{n}$ ;
  • splitting a tokenized document $d$ into $m$ possibly overlapping chunks $d_{i}=t_{i\cdot s}t_{i\cdot s+1}\ldots t_{i\cdot s+w-1}$, where $s$ is a stride and $w$ is a window size. Trailing chunks may have fewer than $w$ tokens.
  • generating $m$ token sequences [CLS] $q$ [SEP] $d_{i}$ [SEP] by concatenating the query with document chunks;
  • processing each sequence with a BERT model to generate a last-layer output for each sequence [CLS] token.
The outcome of this procedure is $m$ [CLS]-vectors $cls_{i}$, which are subsequently aggregated in a model-specific way. Note that PARADE and MaxP/SumP models do not use contextualized embeddings of regular tokens.
BERT MaxP/SumP
These models use a linear layer $F$ to produce $m$ chunk-relevance scores $F(cls_{i})$. Document scores are computed as $\max_{i=1}^{m}F(cls_{i})$ and $\sum_{i=1}^{m}F(cls_{i})$ for MaxP and SumP models, respectively Dai and Callan (2019).
PARADE
These models Li et al. (2024) can be divided into two groups. The first group includes PARADE average, PARADE max, and PARADE attention models, which all use simple approaches to produce an aggregated representation of $m$ [CLS]-vectors $cls_{i}$. To compute a document-relevance score these aggregated representations are passed through a linear layer $F$.
In particular, PARADE average and PARADE max combine $cls_{i}$ using averaging and the element-wise maximum operation, respectively, to generate aggregated representation of $m$ [CLS] tokens $cls_{i}$. [^10] Note that both PARADE average and AvgP vanilla ranker use the same approach to aggregate contextualized embeddings of [CLS] tokens, but they differ in their approach to select document chunks. In particular, AvgP uses non-overlapping chunks while PARADE average relies on the sliding window approach. The PARADE attention model uses a learnable attention Bahdanau et al. (2015) vector $C$ to compute a scalar weight $w_{i}$ of each $cls_{i}$ as follows: $w_{1}w_{2}\ldots w_{m}=softmax(C\cdot cls_{1},C\cdot cls_{2},\ldots,C\cdot cls_{m})$. These weights are used to compute the aggregated representation as $\sum_{i=1}^{m}w_{i}cls_{i}$
PARADE Transformer models aggregates [CLS]-vectors $cls_{i}$ using an additional transformer model $AggregTransf()$, i.e., an aggregator model. The input of the aggregator Transformer is a sequence of $cls_{i}$ vectors prepended with a learnable vector $C$, which plays a role of a [CLS] embedding for $AggregTransf()$. Optionally, one can use learnable positional embedding vectors together with C and $\{cls_{i}\}$. The last-layer representation of the first vector is passed through a linear layer $F$ to produce a relevance score:
$F(AggregTransf(C,cls_{1},cls_{2},\ldots,cls_{m})[0])$ (1)
An aggregator Transformer can be either pretrained or randomly initialized. In the case of a pretrained transformer, we completely discard the embedding layer. Furthermore, if the dimensionality of $cls_{i}$ vectors is different from the dimensionality of input embeddings in $AggregTransf$, we project $cls_{i}$ using a linear transformation.
Miscellaneous models
We attempted to implement additional state-of-the-art models Gao and Callan (2022); Fu et al. (2022). Gao and Callan Gao and Callan (2022) introduced a late-interaction model MORES+, which is a modular long document reranker that uses a sequence-to-sequence transformer in a non-auto-regressive mode. In MORES+ document chunks are first encoded using the encoder-only Transformer model. Then they use a modified decoder Transformer for joint query-to-all-document-chunk cross-attention: This modification changes a causal Transformer into an encoder-only bidirectional Transformer model. As of the moment of writing, the MORES+ model holds the first position on a competitive MS MARCO document leaderboard. [^11] https://microsoft.github.io/MSMARCO-Document-Ranking-Submissions/leaderboard/. However, the authors provide only incomplete implementation which does not fully match the description in the paper (i.e., crucial details are missing). We reimplemented this model to the best of our understanding, but our implementation failed to outperform even BM25.
Inspired by this approach, we managed to implement a late-interaction variant of the PARADE model, which we denoted as PARADE-LATEIR. Similar to the original PARADE model, it splits documents into overlapping chunks. However, it then encodes chunks and queries independently. Next, it uses an interaction Transformer to (1) mix these representations, and (2) combine output using an aggregator Transformer. In total, the model uses three backbone encoder-only Transformers: All of these Transformers are initialized using pretrained models.
Fu et al. (2022) proposed a multi-view interactions-based ranking model (MIR). They implement inter-passage interactions via a multi-view attention mechanism, which enables information propagation at token, sentence, and passage levels. Due to the computational complexity, they restrict these interactions to a set of salient/pivot tokens. However, the paper does not provide enough details regarding the choices of these tokens. There is no software available and authors did not respond to our clarification requests. Thus, this model is also excluded from our evaluation.